### v2

# Function Reconstruction Using Rank-1 Lattices and Lower Sets

Abstract.
Our study focuses on constructing optimal rank-1 lattices that enable exact integration and reconstruction of functions in the Chebyshev space, based on finite index sets. We introduce novel theoretical lower bounds on the minimum number of nodes needed for reconstruction, show equivalence between different plans for reconstruction under certain conditions and propose an innovative algorithm for generating the optimal generator vector of rank-1 lattices. By leveraging the inherent structure of the set of interpolators, our approach ensures admissibility conditions through exhaustive search and verification, outperforming existing methods in terms of computation time and memory usage. Numerical experiments validate the efficiency and practical applicability of our algorithm.
Key words and phrases:
Rank-1 lattices, Function reconstruction, Chebyshev space, Cubature rules, Component-by-component (CBC) strategies, Deterministic algorithms, Approximation theory
2020 Mathematics Subject Classification:
41A10 , 65D30, 41A63, 65D32, 42A15, 65D15
moulay-abdellah.chkifa@polytechnique.org
College of Computing, University Mohammed VI Polytechnic, Morocco. abdelqoddous.moussa@um6p.ma.
1. Introduction
In the field of function integration and reconstruction, rank-1 lattices framework plays a crucial role. Typically, we consider the approximation of the $d$-dimensional integral over the unit cube $[0,1]^{d}$
(1.1) $I_{d}(f)=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}},$
or the approximation of the $d$-variates function ${\boldsymbol{x}}\mapsto f({\boldsymbol{x}})$, by a equal weight rank-1 lattice rules
$Q_{n,{\boldsymbol{z}}}(f)=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{t}}_{i}),\quad\quad\quad\quad{\boldsymbol{x}}\mapsto\sum_{{\boldsymbol{k}}\in\Lambda}Q_{n,{\boldsymbol{z}}}(f\alpha_{\boldsymbol{k}})\alpha_{\boldsymbol{k}}({\boldsymbol{x}})$
where =$\mathcal{L}=\left\{{\boldsymbol{t}}_{i}:=\left\{\frac{i{\boldsymbol{z}}}{n}\right\},i=0,\dots,n-1\right\}\subset[0,1]^{d}$ is the rank-1 lattice generated by ${\boldsymbol{z}}/n$ with ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ and $\{\alpha_{\boldsymbol{k}}\}_{{\boldsymbol{k}}\in\Lambda}$ a basis of multivariate algebraic or trigonometric polynomials. In this article, we are particularly interested in specific spaces of multivariate algebraic or trigonometric polynomials $f$, indexed in the so-called lower sets which are inherent in approximation theory. Our primary focus in this paper is the fast construction of optimal rank-1 lattices, providing through the adequate cubature rules, the exact integration or reconstruction of functions belonging to these spaces. This involves determining the optimal number $n$ of nodes ${\boldsymbol{t}}_{i}$ for the cubature rule and their optimal distribution.
In [cools2010constructing], the authors introduced weighted degrees of exactness and considered anisotropic nodes. To address the exponential growth of required integration points with increasing dimension, they present two algorithms for constructing lattice rules. The first one ensures a prescribed weighted degree of exactness, while the second maximizes the degree at each step. The authors provide theoretical foundations and numerical illustrations to demonstrate the effectiveness of these methods in achieving optimal convergence rates in weighted Korobov spaces [sloan2001].
A key challenge in the domain of rank-1 lattices is constructing the optimal generator vector and determining the appropriate number of nodes (sample size) that satisfy the admissibility conditions for exact integration or reconstruction of functions on a space. The construction of the optimal rank-1 lattice typically employs component-by-component (CBC) strategies. These CBC strategies significantly reduce the search space, providing efficient reconstruction with minimal impact on approximation errors [Kmmerer2014ReconstructingMT]. Rank-1 lattice sampling, coupled with CBC constructions, has gained attention due to its advantageous approximation properties. The efficient use of one-dimensional Fast Fourier Transforms for function reconstruction from sampled points further enhances its appeal [Li_Hickernell_2003].
Recent non-probabilistic algorithms exhibit polynomial runtime concerning both spatial dimension and lattice size [rank1]. Furthermore, [Kammerer2020AFP] presents a linearly complex algorithm for constructing rank-1 lattices that ensures exact integration, demonstrating high efficiency for all functions within a specified truncated Fourier space. Additionally, the authors introduce a probabilistic version of a CBC construction algorithm for rank-1 lattices, improving the computational efficiency of an established deterministic approach. This probabilistic algorithm effectively determines reconstructing rank-1 lattices with high probability, showing efficiency in various scenarios, including the construction of exactly integrating rank-1 lattices.
In [gross2021deterministic], a deterministic algorithm is presented for generating multiple rank-1 lattices, specifically designed for approximating periodic functions with multiple variables. This algorithm decomposes a single rank-1 lattice into smaller ones, streamlining the accurate and effective reconstruction of trigonometric polynomials. The resulting multiple rank-1 lattices are characterized by theoretical limits on sampling points and demonstrate notable computational efficiency.
2. Main Results
This paper addresses the challenge of constructing the optimal generator vector and determining the appropriate number of nodes for the rank-1 lattice both theoretically and numerically.
2.1. Outline
Our paper is organized as follows, in the preliminary section 3, we provide an overview, primarily drawing from [rank1], of the latest advancements in constructing rank-1 lattices and cubature rules for integration and function reconstruction. We focus on a broad finite index set and examine results pertaining to Fourier, cosine, and Chebyshev functional spaces. These findings encompass necessary and sufficient conditions regarding the number and distribution of nodes for exact integration and function reconstruction, alongside theoretical upper bounds on their size. Leveraging the interconnections among the three spaces, known results for integration and reconstruction in Fourier spaces are adapted.
In Section 4, we establish lower bounds on the number of nodes for various approximation plans to achieve exact reconstruction, accompanied by examples illustrating the sharpness of these bounds for each plan. Additionally, we provide theoretical optimal choices of the rank-1 lattice for some useful sets of interpolants.
In Section 5, we introduce the lower set structure and its inherent properties, proving an inequality characterizing lower sets. We also introduce a new parameter of lower sets, namely, extremal elements, which proves useful in providing sufficient conditions to establish links between reconstructions under different plans for useful sets. We demonstrate when these reconstructions are equivalent, which is particularly practical for running algorithms with fewer constraints, as illustrated in Section 6.
Section 7 presents an efficient numerical algorithm for finding near-optimal values of the number of nodes. We offer an in-depth exploration of our algorithm, outlining the steps involved in searching for the optimal choice of the rank-1 lattice parameters. We intricately detail the verification process, emphasizing the pivotal role played by the lower set structure. Furthermore, we validate the efficiency of our algorithm through a complexity analysis of our algorithms in Section 8.1 and experimental results in Section 8.2.2. Finally, our paper concludes in Section 9, summarizing our findings and discussing potential avenues for future research.
2.2. Novelty of the paper
First, in Section 4 we derive new, sharp lower bounds for the minimal number of nodes necessary for exact reconstruction ($n$). These bounds improve upon previous results, such as those provided by [rank1], which focused on upper bounds for $n$. Our findings eliminate a considerable number of non-viable options for the number $n$. Additionally, for standard sets like the simplex sets and the block-type sets, we derive a closed form of the optimal rank-1 lattice. This closed form allows for a better understanding of how the lattice rule depends on the shape of the set, offering deeper insights into the relationship between lattice structures and set geometries.
Second, in Section 6 we demonstrate the equivalence between different reconstruction plans (Plans A, B, and C) under specific conditions for the weighted simplex lattices. To our knowledge, this equivalence is novel and very useful in approximation theory. This result simplifies the implementation of reconstruction algorithms, making them more robust and versatile. By proving these equivalences, we enable practitioners to choose among various reconstruction plans without compromising accuracy or efficiency.
Lastly, in Section 7 we introduce a novel deterministic component-by-component algorithm for constructing nodes of the rank-1 lattice in the Chebyshev setting. Our algorithm is inspired by the work of [gross2021deterministic], who developed a similar algorithm for the Fourier setting. However, our approach is distinct, as it leverages the inherent lower structure of the set of interpolants to optimize our approach, breaking down the problem into more manageable components. By employing exhaustive search techniques and meticulous verification steps, we ensure the fulfillment of admissibility conditions for exact function reconstruction in the truncated space. Unlike [Dick2021ComponentbycomponentCO] and the references therein, who minimize the overall error of approximation within that space, or [Ebert2018ConstructingLP], who construct high-quality quasi-Monte Carlo integration rules for weighted Korobov spaces, our algorithm is valid for every function $f$ independently of its inherent properties [Cools2019FastCC]. This makes our algorithm both fast and memory-efficient, making it suitable for large-scale applications.
3. Preliminaries
We consider the following classical spaces
(3.1) ${\mathcal{F}}^{\rm Four}:=\Big{\{}f\in L^{2}\;|\;f:[0,1]^{d}\mapsto\mathbb{C},f(x)=\sum_{{\boldsymbol{k}}\in\mathbb{Z}^{d}}{\widehat{f}}_{\boldsymbol{k}}e_{\boldsymbol{k}}({\boldsymbol{x}}),\leavevmode\nobreak\ \sum_{{\boldsymbol{k}}\in\mathbb{Z}^{d}}|{\widehat{f}}_{\boldsymbol{k}}|<\infty\Big{\}},$
(3.2) ${\mathcal{F}}^{\rm cos}:=\Big{\{}f\in L^{2}\;|\;f:[0,1]^{d}\mapsto\mathbb{R},f(x)=\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}{\widehat{f}}_{\boldsymbol{k}}\phi_{\boldsymbol{k}}({\boldsymbol{x}}),\leavevmode\nobreak\ \sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}|{\widehat{f}}_{\boldsymbol{k}}|<\infty\Big{\}},$
(3.3) ${\mathcal{F}}^{\rm Cheb}:=\Big{\{}f\in L^{2}\;|\;f:[-1,1]^{d}\mapsto\mathbb{R},f(x)=\sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}{\widehat{f}}_{\boldsymbol{k}}\eta_{\boldsymbol{k}}({\boldsymbol{x}}),\leavevmode\nobreak\ \sum_{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}}|{\widehat{f}}_{\boldsymbol{k}}|<\infty\Big{\}}.$
The underlying measure is the Lebesgue measure $d{\boldsymbol{x}}$ for the first two spaces and the tensor product arcsine measure $\prod_{j=1}^{d}\frac{dx_{j}}{\pi\sqrt{1-x_{j}^{2}}}$ for the third space. For each space, the basis used in the expansion is a complete orthonormal system. They all have a tensor product structure $e_{\boldsymbol{k}}=\otimes_{j=1}^{d}e_{k_{j}}$, $\phi_{\boldsymbol{k}}=\otimes_{j=1}^{d}\phi_{k_{j}}$, $\eta_{\boldsymbol{k}}=\otimes_{j=1}^{d}\eta_{k_{j}}$, where $e_{0}\equiv 1$, $\phi_{0}\equiv 1$, $\eta_{0}\equiv 1$ otherwise $e_{k}(x)=e^{2i\pi kx}$, $\phi_{k}(x)=\sqrt{2}\cos(\pi kx)$ and $\eta_{k}(x)=\sqrt{2}T_{k}(x)$, where $T_{k}$ is the $k$-th Chebyshev polynomial defined by the identity:
$T_{k}(x)=\cos(k\arccos(x)),\quad k=0,1,2,...,\quad\text{for }x\in[-1,1].$
We recall the following
  • $(e_{k})_{k\in\mathbb{Z}}$ is a complete orthonormal system for $L^{2}([0,1],dx)$,
  • $(\phi_{k})_{k\in\mathbb{N}_{0}}$ is a complete orthonormal system for $L^{2}([0,1],dx)$
  • $(\eta_{k})_{k\in\mathbb{N}_{0}}$ is a complete orthonormal system for $L^{2}([-1,1],dx/{[\pi\sqrt{1-x^{2}]}})$.
In our paper, we designate the first space ${\mathcal{F}}^{\mathrm{Four}}$ as the Fourier space, representing periodic functions over the unit cube $[0,1]^{d}$ characterized by absolutely converging Fourier series. The second space, denoted ${\mathcal{F}}^{\mathrm{cos}}$, is termed the cosine space, of periodic functions over $[0,1]^{d}$ with absolutely converging cosine series. The third space, ${\mathcal{F}}^{\mathrm{Cheb}}$, denotes the space of functions possessing absolutely converging Chebyshev series.
Given an arbitrary finite set $\Lambda$ of indices in $\mathbb{Z}^{d}$ (respectively in $\mathbb{N}_{0}^{d}$), the subspace ${\mathcal{F}}_{\Lambda}^{\rm Four}$ (respectively ${\mathcal{F}}_{\Lambda}^{\rm cos}$, ${\mathcal{F}}_{\Lambda}^{\rm Cheb}$) is the space of all functions whose Fourier (respectively cosine, Chebyshev) series is supported solely on $\Lambda$, i.e., $\widehat{f}_{\boldsymbol{k}}=0$ for ${\boldsymbol{k}}\notin\Lambda$. For instance
(3.4) $f\in{\mathcal{F}}_{\Lambda}^{\rm Four}:\quad f(x)=\sum_{{\boldsymbol{k}}\in\Lambda}\widehat{f}_{\boldsymbol{k}}e^{2\pi\mathrm{i}{\boldsymbol{k}}\cdot{\boldsymbol{x}}}.$
3.1. Approximation requirements
For the sake of clarity, in this section we will employ a unified notation for the spaces and bases we have introduced. We use notation $(\alpha_{{\boldsymbol{k}}})_{{\boldsymbol{k}}}$ for the multi-variate complete orthonormal systems and use conjugate notation $\overline{\alpha_{\boldsymbol{k}}}$. The integral $I(f)$ and inner product $\langle f,g\rangle=I(f\overline{g})$ are understood from the context. We will also use ${\mathcal{F}}_{\Lambda}$ for truncated spaces, without specifying Fourier, cosine or Chebyshev.
Placing ourselves in one of the three spaces, a cubature (or quadrature) is simply defined by $Q_{n}(f)=\sum_{j=1}^{n}w_{j}f({\boldsymbol{t}}_{i})$ for given weights $w_{j}>0$ and given nodes ${\boldsymbol{t}}_{i}$ belonging to the underlying domain. One can be interested in the following properties
  • Integral exactness. To demand that the cubature is exact for all functions which are supported solely on $\Lambda$, i.e., $Q_{n}(f)=I(f)$ for all $f\in{\mathcal{F}}_{\Lambda}$. This holds if and only if
    (3.5) $Q_{n}(\alpha_{\boldsymbol{k}})=I(\alpha_{\boldsymbol{k}})=\delta_{{\boldsymbol{k}},{\boldsymbol{0}}},\quad\quad\quad{\boldsymbol{k}}\in\Lambda,$
    where $\delta$ is the Kronecker symbol, i.e. the cubature formula integrates exactly all basis functions $\alpha_{\boldsymbol{k}}$ for ${\boldsymbol{k}}\in\Lambda$.
  • Function reconstruction. To demand that an approximation operator using $Q_{n}$ reconstruct all functions which are supported solely on $\Lambda$. We consider
    (3.6) $f^{a}({\boldsymbol{x}})=\sum_{{\boldsymbol{k}}\in\Lambda}\widehat{f}_{\boldsymbol{k}}^{a}\leavevmode\nobreak\ \alpha_{\boldsymbol{k}}({\boldsymbol{x}}),\quad\quad\quad{\widehat{f}}_{\boldsymbol{k}}^{a}=Q_{n}(f\overline{\alpha_{{\boldsymbol{k}}}}),$
    where each coefficient ${\widehat{f}}_{\boldsymbol{k}}=\langle f,\alpha_{\boldsymbol{k}}\rangle=I(f\overline{\alpha_{\boldsymbol{k}}})$ is replaced by ${\widehat{f}}_{\boldsymbol{k}}^{a}$ obtained by the cubature formula applied to $f\overline{\alpha_{\boldsymbol{k}}}$. We assume the “non-aliasing” condition
    ${\widehat{f}}_{\boldsymbol{k}}^{a}={\widehat{f}}_{\boldsymbol{k}}\quad\quad\quad\forall{\boldsymbol{k}}\in\Lambda\text{ and }f\in\mathcal{F}_{\Lambda},$
    then the operator defined by $f\in{\mathcal{F}}\mapsto f^{a}\in{\mathcal{F}}_{\Lambda}$ is a reconstruction operator (if other coefficients $\widehat{f}_{{\boldsymbol{k}}^{\prime}}$ with ${\boldsymbol{k}}^{\prime}\neq{\boldsymbol{k}}$ contribute to $\widehat{f}_{\boldsymbol{k}}^{a}$ then this is called “aliasing”). It is easily verified that non-aliasing holds if and only if
    (3.7) $Q_{n}(\alpha_{\boldsymbol{k}}\overline{\alpha_{{\boldsymbol{k}}^{\prime}}})=\delta_{{\boldsymbol{k}},{\boldsymbol{k}}^{\prime}}=\langle\alpha_{\boldsymbol{k}},\alpha_{{\boldsymbol{k}}^{\prime}}\rangle,\quad\quad\quad{\boldsymbol{k}},{\boldsymbol{k}}^{\prime}\in\Lambda.$
  • Function reconstruction and bi-orthonormality. As far as the reconstruction using cubatures is concerned, a different approach based on bi-orthonormality is viable. More precisely, we assume given $(\beta_{\boldsymbol{k}})_{{\boldsymbol{k}}\in\Lambda}$ an arbitrary set of functions in ${\mathcal{F}}$, which we assume is bi-orthonormal to $(\alpha_{\boldsymbol{k}})_{{\boldsymbol{k}}\in\Lambda}$, i.e. $\langle\alpha_{\boldsymbol{k}},\beta_{{\boldsymbol{k}}^{\prime}}\rangle=\delta_{{\boldsymbol{k}},{\boldsymbol{k}}^{\prime}}$. We consider the approximation $f^{b}$ by
    (3.8) $f^{b}({\boldsymbol{x}})=\sum_{{\boldsymbol{k}}\in\Lambda}{\widehat{f}}_{\boldsymbol{k}}^{b}\leavevmode\nobreak\ \alpha_{\boldsymbol{k}}({\boldsymbol{x}}),\quad{\widehat{f}}_{\boldsymbol{k}}^{b}=\frac{Q_{n}(f\overline{\beta_{{\boldsymbol{k}}}})}{c_{\boldsymbol{k}}},$
    for some coefficients $c_{\boldsymbol{k}}\neq 0$. In contrast to approximation $f^{a}$, the coefficient ${\widehat{f}}_{\boldsymbol{k}}^{b}$ are not defined using the functions $\alpha_{\boldsymbol{k}}$. Demanding that this scheme reproduce ${\mathcal{F}}_{\Lambda}$, which is also equivalent to reproducing all the $\alpha_{\boldsymbol{k}}$ for ${\boldsymbol{k}}\in\Lambda$, holds if and only if
    (3.9) $Q_{n}(\alpha_{\boldsymbol{k}}\overline{\beta_{{\boldsymbol{k}}^{\prime}}})=c_{\boldsymbol{k}}\delta_{{\boldsymbol{k}},{\boldsymbol{k}}^{\prime}}=c_{\boldsymbol{k}}\langle\alpha_{\boldsymbol{k}},\beta_{{\boldsymbol{k}}^{\prime}}\rangle,\quad\quad\quad{\boldsymbol{k}},{\boldsymbol{k}}^{\prime}\in\Lambda.$
We can cast the above requirement in a duality framework. For integral exactness, one is interested in finding a linear form $\alpha_{\boldsymbol{0}}^{*}:f\mapsto\alpha_{\boldsymbol{0}}^{*}(f)$ over ${\mathcal{F}}$ satisfying $\alpha_{\boldsymbol{0}}^{*}(\alpha_{\boldsymbol{k}})=\delta_{{\boldsymbol{0}},{\boldsymbol{k}}}$ for any ${\boldsymbol{k}}\in\Lambda$, since such form yields $\alpha_{\boldsymbol{0}}(f)=\widehat{f}_{\boldsymbol{0}}=I(f)$ for any $f\in{\mathcal{F}}_{\Lambda}$. As for reconstruction, one is interested in finding a family of linear forms $\alpha_{\boldsymbol{h}}^{*}:f\mapsto\alpha_{\boldsymbol{h}}^{*}(f)$ indexed in $\Lambda$ satisfying $\alpha_{\boldsymbol{h}}^{*}(\alpha_{\boldsymbol{k}})=\delta_{{\boldsymbol{h}},{\boldsymbol{k}}}$ for any ${\boldsymbol{h}},{\boldsymbol{k}}\in\Lambda$, in which case the operator $\sum_{{\boldsymbol{h}}\in\Lambda}\alpha_{\boldsymbol{h}}^{*}(f)\leavevmode\nobreak\ \alpha_{\boldsymbol{h}}$ reconstructs all functions in ${\mathcal{F}}_{\Lambda}$. The forms may or may not depend on $\Lambda$. The linear forms $f\mapsto{\widehat{f}_{\boldsymbol{h}}}^{\square}$ where $\square\in\{a,b,c\}$ introduced later were proposed and investigated on [rank1].
3.2. Rank-1 lattices rules
Rank-1 lattices are studied for integration, reconstruction and approximation in the Fourier space. Given the generating vector ${\boldsymbol{z}}\in\mathbb{Z}^{d}$, the $n$ points of a rank-1 lattice are specified by
(3.10) ${\boldsymbol{t}}_{i}=\frac{i{\boldsymbol{z}}\bmod n}{n}\in[0,1[^{d}\quad\text{for }\quad i=0,\ldots,n-1.$
For a function $f$ in the Fourier (respectively Cosine or Chebyshev) space, the average of the function values at the lattice points (respectively the tent-transformed [Cools2016TenttransformedLR] or tent-transformed cosine-transformed lattice points) are given by
(3.11) $\frac{1}{n}\sum_{i=0}^{n-1}f\left({\boldsymbol{t}}_{i}\right),\quad\frac{1}{n}\sum_{i=0}^{n-1}f\left(\varphi_{\rm tent}({\boldsymbol{t}}_{i})\right),\quad\frac{1}{n}\sum_{i=0}^{n-1}f\left(\cos\left(2\pi{\boldsymbol{t}}_{i}\right)\right).$
[…]
(3.12) $\frac{1}{n}\sum_{i=0}^{n-1}e^{{\mathrm{i}}2\pi i{{\boldsymbol{h}}\cdot{\boldsymbol{z}}}/n}=\begin{cases}1&\text{ if }{\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}0,\\ 0&\text{ otherwise }.\end{cases}$
The character property, and also the property $e_{\boldsymbol{k}}\overline{e_{\boldsymbol{k}}^{\prime}}=e_{{\boldsymbol{k}}-{\boldsymbol{k}}^{\prime}}$, allows to reformulate integral exactness and reconstruction requirements discussed earlier.
Henceforth, we use # to denote the cardinality of a set. For any multiindex ${\boldsymbol{k}}\in\mathbb{Z}^{d}$, we write $|{\boldsymbol{k}}|_{0}:=\#\{1\leq j\leq d:k_{j}\neq 0\}$ to represent the number of nonzero coordinates of ${\boldsymbol{k}}$. For $\sigma\in\{\pm 1\}^{d}$ and ${\boldsymbol{k}}\in\mathbb{Z}^{d}$, we define $\sigma({\boldsymbol{k}}):=(\sigma_{1}k_{1},\ldots,\sigma_{d}k_{d})$, indicating that the sign changes in $\sigma$ are applied component-wise to ${\boldsymbol{k}}$. For any ${\boldsymbol{k}}\in\mathbb{Z}^{d}$, we define
$S_{{\boldsymbol{k}}}:=\left\{\sigma\in\{\pm 1\}^{d}:\sigma_{j}=+1\text{ for each }j\text{ where }k_{j}=0\right\}$
to denote a set of unique sign changes for ${\boldsymbol{k}}$. Clearly, we have #$\#S_{{\boldsymbol{k}}}=2^{|{\boldsymbol{k}}|_{0}}$.
Lemma 3.1 (See [Kmmerer2014ReconstructingMT, Potts2016SparseHF]).
Let $\Lambda\subset\mathbb{Z}^{d}$ be an arbitrary index set. A lattice rank-1 rule $Q_{n}$ with $n$ points and generating vector ${\boldsymbol{z}}$
  • integrates exactly all functions $f\in{\mathcal{F}}^{\rm Four}_{\Lambda}$ iff
    (3.13) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0\quad\mbox{for all}\quad{\boldsymbol{h}}\in\Lambda\setminus\{{\boldsymbol{0}}\},$
  • reconstructs exactly the Fourier coefficients of all $f\in{\mathcal{F}}^{\rm Four}_{\Lambda}$ by ${\widehat{f}^{a}}=Q_{n}(fe_{{\boldsymbol{k}}})$ iff
    (3.14) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda.$
Requirements for integral exactness and reconstruction along the same lines of Lemma 3.1 hold with spaces ${\mathcal{F}}_{\Lambda}^{\rm cos}$, ${\mathcal{F}}_{\Lambda}^{\rm Cheb}$. The lemma combined with identities
(3.15) $\phi_{\boldsymbol{k}}=\frac{1}{2^{|{\boldsymbol{k}}|_{0}}}\sum_{\sigma\in{\mathcal{S}}_{\boldsymbol{k}}}e_{\sigma({\boldsymbol{k}})},\quad\quad\quad\phi_{\boldsymbol{k}}\phi_{{\boldsymbol{k}}^{\prime}}=\frac{1}{2^{|{\boldsymbol{k}}|_{0}+|{\boldsymbol{k}}^{\prime}|_{0}}}\sum_{\begin{subarray}{c}\sigma\in{\mathcal{S}}_{\boldsymbol{k}}\\ \sigma^{\prime}\in{\mathcal{S}}_{{\boldsymbol{k}}^{\prime}}\end{subarray}}e_{\sigma({\boldsymbol{k}})+\sigma^{\prime}({\boldsymbol{k}}^{\prime})},$
which are valid for any ${\boldsymbol{k}},{\boldsymbol{k}}^{\prime}\in\mathbb{N}_{0}^{d}$ yields the results below. The “if” directions in Lemma 3.2 and Lemma 3.3 (plan A) follows. The “only if” directions also hold and were established in [rank1].
Lemma 3.2 (See [rank1]).
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be an arbitrary index set. A tent-transformed lattice rule of a rank-1 lattice rule $Q_{n}$ with $n$ points and generating vector ${\boldsymbol{z}}$ integrates exactly all functions $f\in{\mathcal{F}}^{\rm cos}_{\Lambda}$ iff
(3.16) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0\quad\mbox{for all}\quad{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)\setminus\{{\boldsymbol{0}}\},$
where
${\mathcal{M}}(\Lambda):=\left\{\sigma({\boldsymbol{k}}):{\boldsymbol{k}}\in\Lambda,\quad\sigma\in\{\pm 1\}^{d}\right\}=\bigcup_{{\boldsymbol{k}}\in\Lambda}\{\sigma({\boldsymbol{k}}):\sigma\in S_{{\boldsymbol{k}}}\}.$
In order to present rigorously the reconstruction results established in [rank1], we need to give an account on the cubature introduced their-in. Given a rank-1 lattice rule $Q_{n}$ with $n$ points and generating vector ${\boldsymbol{z}}$, we associate to ${\boldsymbol{k}}\in\mathbb{N}_{0}^{d}$
(3.17) ${\widehat{f}}_{\boldsymbol{k}}^{a}$ $:=Q_{n}\big{(}(f\circ\varphi_{\text{tent }})(\phi_{\boldsymbol{k}}\circ\varphi_{\text{tent }})\big{)},$
(3.18) ${\widehat{f}}_{\boldsymbol{k}}^{b}$ $:=Q_{n}\big{(}(f\circ\varphi_{\text{tent }})\sqrt{2}^{|{\boldsymbol{k}}|_{0}}\cos(2\pi{\boldsymbol{k}}\cdot{\boldsymbol{x}})\big{)},$
(3.19) ${\widehat{f}}_{\boldsymbol{k}}^{c}$ $:=Q_{n}\big{(}(f\circ\varphi_{\text{tent }})\sqrt{2}^{|{\boldsymbol{k}}|_{0}}\cos(2\pi{\boldsymbol{k}}\cdot{\boldsymbol{x}})\big{)}/{c_{{\boldsymbol{k}}}},$
where $c_{{\boldsymbol{k}}}:=\#\{{\sigma}\in{\mathcal{S}}_{\boldsymbol{k}}:\sigma({\boldsymbol{k}})\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{k}}\cdot{\boldsymbol{z}}\}$.
The reconstruction requirement (plan A) is based on ${\widehat{f}}_{\boldsymbol{k}}^{a}$ and is relatively straightforward. The requirements can be weakened employing the bi-orthonormality trick. By observing that
[…]
one can use functions $\beta_{\boldsymbol{k}}=\sqrt{2}^{|{\boldsymbol{k}}|_{0}}\cos(2\pi{\boldsymbol{k}}\cdot{\boldsymbol{x}})$ in order to carry the bi-orthonormality trick, which we have briefly discussed. This observation along the requirements (plan B/C) below where also derived and thoroughly discussed in [rank1].
Lemma 3.3 ([rank1]).
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be an arbitrary index set. A tent-transformed lattice rule $Q_{n}$ of a rank-1 lattice rule $Q_{n}$ with $n$ points and generating vector ${\boldsymbol{z}}$ reconstructs exactly the cosine coefficients of all $f\in{\mathcal{F}}^{\rm cos}_{\Lambda}$
  • (plan A) by ${\widehat{f}}={\widehat{f}^{a}}$ iff
    (3.21) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda),$
  • (plan B) by ${\widehat{f}}={\widehat{f}^{b}}$ iff
    (3.22) $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\leavevmode\nobreak\ \sigma({\boldsymbol{h}})\neq{\boldsymbol{h}}^{\prime},$
  • (plan C) by ${\widehat{f}}={\widehat{f}^{c}}$ iff
    (3.23) $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime},$
Let us discuss the difference between requirements (3.21), (3.22) and (3.23). First, it is obvious that (3.21) is stronger than (3.22). Since $\Lambda\subset\mathbb{N}_{0}^{d}$, then for any ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda$ and $\sigma\in{\mathcal{S}}_{h}$ one clearly has $\sigma({\boldsymbol{h}})\neq{\boldsymbol{h}}^{\prime}$ except if ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}$ and $\sigma=\rm id$. Requiring (3.22) is therefore equivalent to requiring (3.23) and in addition that $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\in\Lambda$ and $\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}-\{\rm id\}$. In other words, in (3.23) it is allowed that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}$ occur in the set $\{\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\}_{\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}-\{\rm id\}}$.
For latter use, we introduce the proper notation for admissibility with plan A, B and C. Given $\Lambda\subset\mathbb{N}_{0}^{d}$ an arbitrary set of indices, we let ${\mathcal{P}}_{0}(\Lambda)$ be the set of couples $(n,{\boldsymbol{z}})\in\mathbb{N}\times\mathbb{Z}^{d}$ satisfying (3.16). Similarly, we introduce ${\mathcal{P}}_{A}(\Lambda)$, ${\mathcal{P}}_{B}(\Lambda)$ and ${\mathcal{P}}_{C}(\Lambda)$ to be the sets of couples $(n,{\boldsymbol{z}})\in\mathbb{N}\times\mathbb{Z}^{d}$ satisfying (3.21), (3.22), and (3.23) respectively. We say that $(n,{\boldsymbol{z}})$ is admissible for plan $\square$ if $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{\square}(\Lambda)$.
The discussion above shows that ${\mathcal{P}}_{A}(\Lambda)\subset{\mathcal{P}}_{B}(\Lambda)\subset{\mathcal{P}}_{C}(\Lambda)$. For the sake of completeness, we summarize in the following proposition.
Proposition 3.4.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ finite, the following implications hold
$(n,{\boldsymbol{z}})$ is admissible for plan A $\implies$ $(n,{\boldsymbol{z}})$ is admissible for plan B $\implies$ $(n,{\boldsymbol{z}})$ is admissible for plan C.
Remark 3.5.
We denote ${\mathcal{S}}_{{\boldsymbol{h}}}^{*}={\mathcal{S}}_{{\boldsymbol{h}}}-\{\rm id\}$. We remark that ${\mathcal{S}}_{{\boldsymbol{0}}}=\{\rm id\}$ and therefore ${\mathcal{S}}_{{\boldsymbol{0}}}^{*}=\emptyset$. Since $\Lambda\subset\mathbb{N}_{0}^{d}$, one can easily verify that the extra requirement on admissibility condition of every plan with respect to others is as outlined below. By O-O, we simply mean ${\boldsymbol{h}}\in\Lambda\mapsto{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ is one-to-one.
$\begin{array}[]{|l l|c|c|c|c|}\hline\cr\text{Requirement}&&\text{O-O}&\text{C}&\text{ B}&\text{A}\\ \hline\cr{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}&{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda&\checkmark&\checkmark&\checkmark&\checkmark\\ {\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}&{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}^{*}&&\checkmark&\checkmark&\checkmark\\ {\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ \not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}&{\boldsymbol{h}}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}^{*}&&&\checkmark&\checkmark\\ \sigma^{\prime}({\boldsymbol{h}}^{\prime})\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}&{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}^{*},\leavevmode\nobreak\ \sigma^{\prime}\in{\mathcal{S}}_{{\boldsymbol{h}}^{\prime}}^{*}.&&&&\checkmark\\ \hline\cr\end{array}$
For any given ${\boldsymbol{z}}\in\mathbb{Z}^{d}$, we let $n^{*}_{0}(\Lambda,{\boldsymbol{z}})$ be the smallest possible integer $n$ s.t. $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{0}(\Lambda)$. By convention $n^{*}_{0}(\Lambda,{\boldsymbol{z}})=+\infty$ if no such $n$ exists, e.g. the case ${\boldsymbol{z}}={\boldsymbol{0}}$, and $\lambda\neq\{0\}$. We define similarly the integers $n^{*}_{A}(\Lambda,{\boldsymbol{z}})$, $n^{*}_{B}(\Lambda,{\boldsymbol{z}})$ and $n^{*}_{C}(\Lambda,{\boldsymbol{z}})$.
We introduce a dedicated notation $n^{*}_{0}(\Lambda)$ for the smallest possible integer $n$ such that $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{0}(\Lambda)$ for some ${\boldsymbol{z}}$. We define similarly the notation $n^{*}_{A}(\Lambda)$, $n^{*}_{B}(\Lambda)$ and $n^{*}_{C}(\Lambda)$. Obviously $n^{*}_{\square}(\Lambda)\geq 2$ where $\square$ is 00 or A/B/C. Upper bounds $n_{\square}^{*}(\Lambda)\leq p_{\square}^{*}(\Lambda)$ are provided in [rank1]. Namely
(3.24) $\begin{array}[]{cl}p_{0}^{*}(\Lambda)&\text{smallest prime }\geq\max\left\{{\#({\mathcal{M}}(\Lambda)\backslash\{0\}}+2,2q\right\}/2,\\ p_{A}^{*}(\Lambda)&\text{smallest prime }\geq\max\left\{{\#({\mathcal{M}}(\Lambda)\oplus{\mathcal{M}}(\Lambda))}+2,4q\right\}/2,\\ p_{B}^{*}(\Lambda)&\text{smallest prime }\geq\max\left\{{\#(\Lambda\oplus{\mathcal{M}}(\Lambda))},2q\right\},\\ p_{C}^{*}(\Lambda)&\text{smallest prime }\geq\max\left\{{\#(\Lambda)\#({\mathcal{M}}(\Lambda))},2q\right\},\end{array}$
where $q=\max\{\|{\boldsymbol{k}}\|_{\infty}:{\boldsymbol{k}}\in\Lambda\}$.
In the next section, we derive new lower bounds on $n$, the number of rank-$1$ lattice points needed for exact integration and the exact reconstruction using plan A, plan B, or plan C.
4. Necessary condition for the exact reconstruction
4.1. Necessary condition for plans A/B/C
In this section, we derive lower bounds on the numbers $n^{*}_{A}(\Lambda)$, $n^{*}_{B}(\Lambda)$ and $n^{*}_{C}(\Lambda)$ of minimal rank-$1$ lattice points needed for exact reconstruction over ${\mathcal{F}}^{\rm cos}_{\Lambda}$ using plan A, plan B, and plan C respectively.
Proposition 4.1.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be an arbitrary index set of finite cardinality.
  • for any value $n<\#({\mathcal{M}}({\Lambda}))$, there exists no ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ s.t.
    (4.1) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{ for all }\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda).$
    If ${\boldsymbol{0}}\not\in\Lambda$, the conclusion also holds for $n=\#({\mathcal{M}}({\Lambda}))$.
  • for any value $n<\#(\{{\boldsymbol{h}},-{\boldsymbol{h}}\}_{{\boldsymbol{h}}\in\Lambda})$, there exists no ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ s.t.
    (4.2) $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\leavevmode\nobreak\ \sigma({\boldsymbol{h}})\neq{\boldsymbol{h}}^{\prime},$
    and if ${\boldsymbol{0}}\not\in\Lambda$, the conclusion also holds for $n=\#(\{{\boldsymbol{h}},-{\boldsymbol{h}}\}_{{\boldsymbol{h}}\in\Lambda})$,
  • for any value $n<\#(\Lambda)$, there exists no ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ s.t.
    (4.3) $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{for all}\quad{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}.$
Proof.
We consider ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ arbitrary and the associated map
${\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)\mapsto{\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n\in\{0,\dots,n-1\}.$
If $n<\#({\mathcal{M}}({\Lambda}))$, then by the pigeonhole principle the map cannot be injective, which proves the first claim. We assume now $n=\#({\mathcal{M}}({\Lambda}))$ but ${\boldsymbol{0}}\not\in\Lambda$. Two situations are to be considered. If the image of ${\mathcal{M}}({\Lambda})$ is strictly included in $\{0,\dots,n-1\}$, then again by the pigeonhole principle the map cannot be injective. Otherwise, all the remainders $10,1,\cdots,n-$ 1 are obtained by the mapping. In particular, there exists ${\boldsymbol{h}}\in{\mathcal{M}}({\Lambda})$ satisfying ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}0$. This implies that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ which again make the map not injective since $-{\boldsymbol{h}}\in{\mathcal{M}}({\Lambda})$ and ${\boldsymbol{h}}\neq-{\boldsymbol{h}}$. The requirement (4.2) implies that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda$ (by letting $\sigma=id$) and that $-{\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda$ (by letting $\sigma=-id$ and checking that $-{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}$ holds except if ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}={\boldsymbol{0}}$). Also, since $x\equiv_{n}y$ if and only if $-x\equiv_{n}-y$, the requirement implies $-{\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}-{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda$. We introduce the set $\Lambda^{\pm}=\{{\boldsymbol{h}},-{\boldsymbol{h}}\}_{{\boldsymbol{h}}\in\Lambda}$. The requirement thus implies that the map ${\boldsymbol{h}}\in\Lambda^{\pm}\mapsto({\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ is injective. But, this cannot hold if $n<\#(\Lambda^{\pm})$. If ${\boldsymbol{0}}\not\in\Lambda$, the same argument used above applies since as for ${\mathcal{M}}(\Lambda)$, the set $\Lambda^{\pm}$ satisfies ${\boldsymbol{h}}\in\Lambda^{\pm}\Longrightarrow-{\boldsymbol{h}}\in\Lambda^{\pm}$. Requirement (4.3) implies clearly that the map ${\boldsymbol{h}}\in\Lambda\mapsto({\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ is injective (by letting $\sigma=id$). We cannot have this if $n<\#(\Lambda)$. ∎
We note that #$\#(\Lambda^{\pm})$ is equal to $2\#(\Lambda)$ if ${\boldsymbol{0}}\not\in\Lambda$ and to $12\#(\Lambda)-$ 1 otherwise. The lemma provides lower bounds on the number of points that are necessary to perform exact reconstruction over ${\mathcal{F}}^{\rm cos}_{\Lambda}$ using plan A, plan B, and plan C. In particular, we have proved that for any finite set $\Lambda\subset\mathbb{N}_{0}^{d}$
(4.4) $\begin{array}[]{l}n^{*}_{A}(\Lambda)\geq l^{*}_{A}(\Lambda):=\#({\mathcal{M}}({\Lambda}))+1_{{\boldsymbol{0}}\not\in\Lambda},\\ n^{*}_{B}(\Lambda)\geq l^{*}_{B}(\Lambda):=2\#(\Lambda)+\delta_{\Lambda,{\boldsymbol{0}}},\\ n^{*}_{C}(\Lambda)\geq l^{*}_{C}(\Lambda):=\#(\Lambda),\end{array}$
where
$\delta_{\Lambda,{\boldsymbol{0}}}=1$ if ${\boldsymbol{0}}\not\in\Lambda$ and $\delta_{\Lambda,{\boldsymbol{0}}}=-1$ if ${\boldsymbol{0}}\in\Lambda$.
Next, we demonstrate that $n^{*}_{A}(\Lambda)=\#({\mathcal{M}}({\Lambda}))$ is optimal in some cases, in particular in any dimension for $\Lambda$ of “block” type. It is worth noting that while the other lower bounds are not optimal, they provide a favorable initial guess to initiate algorithms.
(a) block
(b) cross
(c) simplex
Figure 1. Sets ${\mathcal{B}}_{{\boldsymbol{k}}}$, ${\mathcal{C}}_{{\boldsymbol{k}}}$ and ${\mathcal{S}}_{{\boldsymbol{1}},k}$ and associated mirrored sets in dimension 2.
4.2. Sharpness of the necessary condition
For ${\boldsymbol{k}},{\boldsymbol{k}}^{\prime}\in\mathbb{Z}^{d}$, ${\boldsymbol{k}}^{\prime}\leq{\boldsymbol{k}}$ is to be interpreted component-wise, i.e., $k^{\prime}_{j}\leq k_{j}$ for all $j$. We introduce the notation ${\mathcal{B}}_{{\boldsymbol{k}}}$, ${\mathcal{C}}_{{\boldsymbol{k}}}$ and ${\mathcal{S}}_{{\boldsymbol{1}},k}$ for ${\mathcal{B}}_{{\boldsymbol{k}}}:=\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{2}:{\boldsymbol{h}}\leq{\boldsymbol{k}}\}$, ${\mathcal{C}}_{{\boldsymbol{k}}}:=\{{\boldsymbol{h}}\in{\mathcal{B}}_{{\boldsymbol{k}}}:h_{1}h_{2}=0\}$ where ${\boldsymbol{k}}\in\mathbb{N}_{0}^{2}$ and ${\mathcal{S}}_{{\boldsymbol{1}},k}=\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{2}:h_{1}+h_{2}\leq k\}$ where $k\in\mathbb{N}_{0}$. Mirrored sets are ${\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}})=\{{\boldsymbol{h}}\in\mathbb{Z}^{2}:-{\boldsymbol{k}}\leq{\boldsymbol{h}}\leq{\boldsymbol{k}}\}$, ${\mathcal{M}}({\mathcal{C}}_{{\boldsymbol{k}}})=\{{\boldsymbol{h}}\in{\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}):h_{1}h_{2}=0\}$, and ${\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k})=\{{\boldsymbol{h}}\in\mathbb{Z}^{2}:|h_{1}|+|h_{2}|\leq k\}$. Cardinalities of all these sets are easily verified
(4.5) #$\begin{array}[]{ll}\#({\mathcal{B}}_{{\boldsymbol{k}}})=(k_{1}+1)(k_{2}+1),&\quad\#({\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}))=(2k_{1}+1)(2k_{2}+1),\\ \#({\mathcal{C}}_{{\boldsymbol{k}}})=k_{1}+k_{2}+1,&\quad\#({\mathcal{M}}({\mathcal{C}}_{{\boldsymbol{k}}}))=2(k_{1}+k_{2})+1,\\ \#({\mathcal{S}}_{{\boldsymbol{1}},k})={(k+1)(k+2)}/2,&\quad\#({\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k}))=2k^{2}+2k+1.\end{array}$
Proposition 4.2.
Let ${\boldsymbol{k}}\in\mathbb{N}_{0}^{2}$, $n=\#({\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}))$ and ${\boldsymbol{z}}=(1,2k_{1}+1)$. Then
(4.6) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{ for all }\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}).$
Proof.
We let ${\boldsymbol{h}}$, ${\boldsymbol{h}}^{\prime}$ be in ${\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}})$ satisfying ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$. We have that $(h_{1}^{\prime}-h_{1})+(h_{2}^{\prime}-h_{2})(2k_{1}+1)=nq$ for some integer $q$. Since $n|q|\leq 2k_{1}+2k_{2}\times(2k_{1}+1)=n-1$, necessarily $q=0$. This in turn implies $(2k_{1}+1)(h_{2}-h_{2}^{\prime})=h_{1}^{\prime}-h_{1}\in\{-k_{1},\dots,k_{1}\}$ which holds if and only if $h_{2}^{\prime}-h_{2}=0$ and $h_{1}^{\prime}-h_{1}=0$, hence ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}$. The proof is complete. ∎
Proposition 4.3.
Let $k\in\mathbb{N}$, $n=\#({\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k}))$ and ${\boldsymbol{z}}=(1,2k+1)$. Then
(4.7) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{ for all }\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k}).$
Proof.
We let ${\boldsymbol{h}}$, ${\boldsymbol{h}}^{\prime}$ be in ${\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k})$ satisfying ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$. We have that $(h_{1}^{\prime}-h_{1})+(h_{2}^{\prime}-h_{2})(2k+1)=nq$ for some integer $q$, which we can assume $\geq 0$ up to permute the roles of ${\boldsymbol{h}}$ and ${\boldsymbol{h}}^{\prime}$. We have $(h_{1}^{\prime}+h_{2}^{\prime})-(h_{1}+h_{2})+2k(h_{2}^{\prime}-h_{2})=nq$, thus $0\leq nq\leq k+k+2k\times 2k=4k^{2}+2k$. In view of $n=2k^{2}+2k+1$, we deduce that $q\in\{0,1\}$. If $q=1$, we reformulate as $(h_{1}^{\prime}-h_{1}+k)+(h_{2}^{\prime}-h_{2})(2k+1)=(k+1)(2k+1)$, hence $(h_{1}^{\prime}-h_{1}+k)$ is a multiple of $(2k+1)$. Since $|h_{1}|,|h_{1}^{\prime}|\leq k$, then necessarily $h_{1}^{\prime}-h_{1}+k=0$ or $h_{1}^{\prime}-h_{1}+k=2k+1$, to which correspond $h_{2}^{\prime}-h_{2}=k+1$ and $h_{2}^{\prime}-h_{2}=-k$ respectively. In the first case, necessarily $h_{1}\geq 0$ and $h_{2}\leq-1$, implying
$k\geq\|{\boldsymbol{h}}\|_{\ell_{1}}=h_{1}-h_{2}=h_{1}^{\prime}+k+k+1-h_{2}^{\prime}\geq 2k+1-\|{\boldsymbol{h}}^{\prime}\|_{\ell_{1}}>k.$
The second case is similar, simply the roles of $h_{1},h_{1}^{\prime}$ and $h_{2},h_{2}^{\prime}$ as swapped. We deduce that necessarily $q=0$. This in turn implies $(2k+1)(h_{2}^{\prime}-h_{2})=(h_{1}-h_{1}^{\prime})\in\{-k,\dots,k\}$ which holds if and only if $h_{2}^{\prime}-h_{2}=0$ and $h_{1}^{\prime}-h_{1}=0$, hence ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}$. The proof is complete. ∎
Remark 4.4.
With ${\boldsymbol{z}}=(k,k+1)$, the same result as above holds. In addition, the proof is much simpler as we quickly derive that $q=0$ and conclude. This is the lattice rule for the Padua points (see [Padua]).
Proposition 4.5.
Let ${\boldsymbol{k}}\in\mathbb{N}^{2}$, $n=\#({\mathcal{B}}_{\boldsymbol{k}})+1$ and ${\boldsymbol{z}}=(1,k_{1}+1)$. Then
(4.8) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\quad\mbox{ for all }\quad{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{C}}_{{\boldsymbol{k}}}).$
Moreover, for any $n\leq\#({\mathcal{B}}_{\boldsymbol{k}})$ and any ${\boldsymbol{z}}$, the above cannot hold.
Proof.
We let ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}$ be in ${\mathcal{M}}({\mathcal{C}}_{{\boldsymbol{k}}})$ satisfying ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$. Then $(h_{1}-h_{1}^{\prime})+(h_{2}-h_{2}^{\prime})(k_{1}+1)=nq$ for some integer $q$, which we assume $\geq 0$. We then use $0\leq nq\leq|h_{1}-h_{1}^{\prime}|+|h_{2}-h_{2}^{\prime}|(\leavevmode\nobreak\ k_{1}+1)$, $n=(k_{1}+1)(k_{2}+1)+1$, and distinguish between the different cases:
  • if $h_{1}=h_{1}^{\prime}=0$: then $nq\leq 2k_{2}(k_{1}+1)<2n$ implying $q=0$ or $q=1$,
  • if $h_{2}=h_{2}^{\prime}=0$: then $nq\leq 2k_{1}<n$ implying $q=0$,
  • if $h_{1}=h_{2}^{\prime}=0$: then $nq\leq k_{1}+k_{2}(k_{1}+1)<n$ implying $q=0$,
  • if $h_{1}^{\prime}=h_{2}=0$: then $nq\leq k_{1}+k_{2}(k_{1}+1)<n$ implying $q=0$.
In the first case, $q=1$ implies $(h_{2}-h_{2}^{\prime}-(k_{2}+1))(k_{1}+1)=1$ but this cannot hold. Therefore, $q=0$ in all cases. In the first and second case, this clearly yields ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}={\boldsymbol{0}}$. In the third case, it implies $-h_{1}^{\prime}+h_{2}(k_{1}+1)=0$ hence in view of $|h_{1}^{\prime}|\leq k_{1}$ necessarily $h_{1}^{\prime}=h_{2}^{\prime}=0$, so that ${\boldsymbol{h}}={\boldsymbol{h}}^{\prime}={\boldsymbol{0}}$. The same holds for the forth case. The first part of the proof is complete. Now we prove that for $n\leq\#({\mathcal{B}}_{\boldsymbol{k}})$ and any ${\boldsymbol{z}}\in\mathbb{Z}^{2}$, it is impossible to have (4.8). For ${\boldsymbol{z}}=(z_{1},z_{2})$ arbitrary, we consider the set
${\mathcal{S}}=\{{\boldsymbol{h}}\cdot{\boldsymbol{z}}{\leavevmode\nobreak\ \text{mod}\leavevmode\nobreak\ }n\}_{{\boldsymbol{h}}\in{\mathcal{B}}_{{\boldsymbol{k}}}}.$
If $n<\#({\mathcal{B}}_{{\boldsymbol{k}}})$, by the pigeonhole principle there must exist ${\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in{\mathcal{B}}_{{\boldsymbol{k}}}$ s.t. ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}$. This in turn implies ${\boldsymbol{l}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{l}}\cdot{\boldsymbol{z}}$ where ${\boldsymbol{l}}^{\prime}=(h_{1}^{\prime}-h_{1},0)$ and ${\boldsymbol{l}}=(0,h_{2}-h_{2}^{\prime})$, which are distinct and both belong to ${\mathcal{M}}({\mathcal{C}}_{{\boldsymbol{k}}})$. If $n=\#({\mathcal{B}}_{{\boldsymbol{k}}})$, then either we have a similar situation otherwise all the remainder $10,\dots,n-$ 1 are taken by $({\boldsymbol{h}}\cdot{\boldsymbol{z}}{\leavevmode\nobreak\ \text{mod}\leavevmode\nobreak\ }n)$ for ${\boldsymbol{h}}$ varying in ${\mathcal{B}}_{{\boldsymbol{k}}}$. In this case, there exists $(a,b)\in{\mathcal{B}}_{{\boldsymbol{k}}}$ s.t. $az_{1}+bz_{2}\equiv_{n}1$. Also, there exist ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{B}}_{{\boldsymbol{k}}}$ s.t. $h_{1}z_{1}+h_{2}z_{2}\equiv_{n}(k_{1}+1)z_{1}$ and $h_{1}^{\prime}z_{1}+h_{2}^{\prime}z_{2}\equiv_{n}(k_{2}+1)z_{2}$, which we reformulate as $h_{2}z_{2}\equiv_{n}(k_{1}+1-h_{1})z_{1}$ and $h_{1}^{\prime}z_{1}\equiv_{n}(k_{2}+1-h_{2}^{\prime})z_{2}$. In particular, if $1\leq h_{1}\leq k_{1}$ we obtain ${\boldsymbol{l}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{l}}^{\prime}\cdot{\boldsymbol{z}}$ where ${\boldsymbol{l}}=(k_{1}+1-h_{1},0)$ and ${\boldsymbol{l}}^{\prime}=(0,h_{2})$, and if $1\leq h_{2}^{\prime}\leq k_{2}$, we obtain ${\boldsymbol{l}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{l}}^{\prime}\cdot{\boldsymbol{z}}$ where ${\boldsymbol{l}}=(h_{1}^{\prime},0)$ and ${\boldsymbol{l}}^{\prime}=(0,k_{2}+1-h_{2}^{\prime})$, with ${\boldsymbol{l}}\neq{\boldsymbol{l}}^{\prime}\in\mathcal{M}({\mathcal{C}}_{{\boldsymbol{k}}})$ in both cases, which yields to the violation of equation (4.8). It only remains the case $h_{1}=h_{2}^{\prime}=0$, which implies
$\left\{\begin{array}[]{rl}h_{2}z_{2}\equiv_{n}(k_{1}+1)z_{1},\\ h_{1}^{\prime}z_{1}\equiv_{n}(k_{2}+1)z_{2}.\end{array}\right.$
Multiplying these equations by $(k_{2}+1)$ and $(k_{1}+1)$ respectively yields
$\left\{\begin{array}[]{rl}h_{2}h_{1}^{\prime}z_{2}\equiv_{n}nz_{1}\equiv_{n}0,\\ h_{2}h_{1}^{\prime}z_{1}\equiv_{n}nz_{2}\equiv_{n}0.\end{array}\right.$
Combining with the equality $az_{1}+bz_{2}\equiv_{n}1$, we obtain $h_{2}h_{1}^{\prime}\equiv_{n}h_{2}h_{1}^{\prime}(az_{1}+bz_{2})\equiv_{n}0$, which implies $n$ divides $h_{2}h_{1}^{\prime}$. In view of $n=(k_{1}+1)(k_{2}+1)$ and $|h_{2}h_{1}^{\prime}|\leq k_{2}k_{1}$ this is only possible if $h_{2}=0$ or $h_{1}^{\prime}=0$. We again obtain ${\boldsymbol{l}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{l}}\cdot{\boldsymbol{z}}$ with ${\boldsymbol{l}}=(-1,0)$ and ${\boldsymbol{l}}^{\prime}=(k_{1},0)$ or ${\boldsymbol{l}}=(0,-1)$ and ${\boldsymbol{l}}^{\prime}=(0,k_{2})$ respectively, which in both cases satisfy ${\boldsymbol{l}}\neq{\boldsymbol{l}}^{\prime}\in\mathcal{M}({\mathcal{C}}_{{\boldsymbol{k}}})$. We conclude that (4.8) cannot hold. The proof is complete. ∎
In each of the three previous propositions, $n$ is the smallest possible number for which the conditions hold. In particular, in dimension $d=2$ one has
(4.9) $\begin{array}[]{rll}n_{A}^{*}({\mathcal{B}}_{\boldsymbol{k}})&=&(2k_{1}+1)(2k_{2}+1),\\ n_{A}^{*}({\mathcal{C}}_{\boldsymbol{k}})&=&(k_{1}+1)(k_{2}+1)+1,\\ n_{A}^{*}({\mathcal{S}}_{{\boldsymbol{1}},k})&=&2k^{2}+2k+1.\end{array}$
The first result also holds in higher dimensions. We now use notation ${\mathcal{B}}_{{\boldsymbol{k}}}:=\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{d}:{\boldsymbol{h}}\leq{\boldsymbol{k}}\}$ where ${\boldsymbol{h}}\in\mathbb{N}_{0}^{d}$ to denote a block in dimension $d$.
Lemma 4.6.
Consider ${\mathcal{B}}_{{\boldsymbol{k}}}$ and let $n_{\boldsymbol{k}}=\#({\mathcal{B}}_{{\boldsymbol{k}}})$. The vector ${\boldsymbol{z}}=(z_{1},\cdots,z_{d})$ defined by $z_{1}=1$ and $z_{i}=\prod_{j=1}^{i-1}(k_{j}+1)$ for $i=2,\dots,d$, yields
[…]
Moreover, $|{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}|<n_{{\boldsymbol{k}}}$ for any ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{B}}_{{\boldsymbol{k}}}$.
Proof.
For the sake of clarity, we distinguish the cases $d=1$, $d=2$ and $d\geq 3$. This also allows us to build up the intuition that led to the particular choice of ${\boldsymbol{z}}$.
$\bullet$ Case $d=1$: ${\mathcal{B}}_{k}=\{0,\dots,k\}$ and $n_{k}=k+1$. Every integer $z_{1}\neq 0$ ensures the map is injective but only $z_{1}=\pm 1$ ensure the second conclusion.
$\bullet$ Case $d=2$: ${\mathcal{B}}_{{\boldsymbol{k}}}=\{0,\dots,k_{1}\}\times\{0,\dots,k_{2}\}$ and $n_{\boldsymbol{k}}=(k_{1}+1)(k_{2}+1)$. We have ${\boldsymbol{h}}\cdot{\boldsymbol{z}}=h_{1}z_{1}+h_{2}z_{2}$. For $h_{2}$ varying in $\{0,\dots,k_{2}\}$, the values $h_{2}z_{2}$ make jumps of length $|z_{2}|$. By setting $z_{1}=1$, so that $h_{1}z_{1}=h_{1}$ takes the contiguous values $0,\dots,k_{1}$, it is sufficient to set $z_{2}=(k_{1}+1)$ in order to ensure ${\boldsymbol{h}}\to{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ is injective over ${\mathcal{B}}_{{\boldsymbol{k}}}$. Moreover $|{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}|\leq|h^{\prime}_{1}-h_{1}|+(k_{1}+1)|h_{2}^{\prime}-h_{2}|\leq k_{1}+(k_{1}+1)k_{2}=n_{\boldsymbol{k}}-1$ for any ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{B}}_{{\boldsymbol{k}}}$.
$\bullet$ Case $d\geq 3$: we have ${\mathcal{B}}_{{\boldsymbol{k}}}=\{0,\dots,k_{1}\}\times{\mathcal{B}}_{\widehat{\boldsymbol{k}}}$, and $n_{{\boldsymbol{k}}}=(k_{1}+1)n_{\widehat{\boldsymbol{k}}}$, where $\widehat{\boldsymbol{k}}=(k_{2},\dots,k_{d})$. In particular, $n_{{\boldsymbol{k}}}-1=k_{1}+(k_{1}+1)(n_{\widehat{\boldsymbol{k}}}-1)$. The idea is to exploit the above properties and set ${\boldsymbol{z}}$ to a choice that allow us to unroll an induction and derive the desired conclusions. We set ${\boldsymbol{z}}=(1,(k_{1}+1)\widehat{\boldsymbol{z}})$, then
${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}=(h_{1}^{\prime}-h_{1})+(k_{1}+1)(\widehat{\boldsymbol{h}}^{\prime}\cdot\widehat{\boldsymbol{z}}-\widehat{\boldsymbol{h}}\cdot\widehat{\boldsymbol{z}}).$
For ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{B}}_{{\boldsymbol{k}}}$, $|h_{1}^{\prime}-h_{1}|\leq k_{1}$. If in addition we assume $|\widehat{\boldsymbol{h}}^{\prime}\cdot\widehat{\boldsymbol{z}}-\widehat{\boldsymbol{h}}\cdot\widehat{\boldsymbol{z}}|<n_{\widehat{\boldsymbol{k}}}$, we obtain $|{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}|\leq k_{1}+(k_{1}+1)(n_{\widehat{\boldsymbol{k}}}-1)=n_{{\boldsymbol{k}}}-1$. Moreover if ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}={\boldsymbol{h}}\cdot{\boldsymbol{z}}$ then
$(h_{1}^{\prime}-h_{1})+(2k_{1}+1)(\widehat{\boldsymbol{h}}^{\prime}\cdot\widehat{\boldsymbol{z}}-\widehat{\boldsymbol{h}}\cdot\widehat{\boldsymbol{z}})=0,$
which in view of $|h_{1}^{\prime}-h_{1}|\leq k_{1}$ implies $h_{1}^{\prime}=h_{1}$ and $\widehat{\boldsymbol{h}}^{\prime}\cdot\widehat{\boldsymbol{z}}=\widehat{\boldsymbol{h}}\cdot\widehat{\boldsymbol{z}}$. Defining $\widehat{\boldsymbol{z}}$ from $\widehat{\boldsymbol{k}}$ in a similar fashion to ${\boldsymbol{z}}$ from ${\boldsymbol{k}}$ allows us to re-iterate the argument and yield $h_{2}^{\prime}=h_{2},\dots,\leavevmode\nobreak\ h_{d}^{\prime}=h_{d}$. The choice ${\boldsymbol{z}}$ suggested in the lemma satisfy the desired properties. The proof is complete. ∎
The lemma is not exclusive to blocks anchored to ${\boldsymbol{0}}$. Over all translated block ${\boldsymbol{a}}+{\mathcal{B}}_{\boldsymbol{k}}$ for ${\boldsymbol{a}}\in\mathbb{N}_{0}^{d}$, the same dependence of ${\boldsymbol{z}}$ in ${\boldsymbol{k}}$ yields the exact same conclusions. In particular, since ${\mathcal{M}}({\mathcal{B}}_{\boldsymbol{k}})=-{\boldsymbol{k}}+{\mathcal{B}}_{2{\boldsymbol{k}}}$, we have the following.
Lemma 4.7.
Consider ${\mathcal{B}}_{{\boldsymbol{k}}}$ and let $m_{\boldsymbol{k}}=\#({\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}))$. The vector ${\boldsymbol{z}}=(z_{1},\cdots,z_{d})$ defined by $z_{1}=1$ and $z_{i}=\prod_{j=1}^{i-1}(2k_{j}+1)$ for $i=2,\dots,d$, yields
[…]
Moreover, $|{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}|<m_{{\boldsymbol{k}}}$ for any ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}})$.
In the previous lemmas, many others choices of ${\boldsymbol{z}}$ yield the same conclusions. By building ${\boldsymbol{z}}$ as suggested but in some order $z_{\pi(1)},\dots,z_{\pi(d)}$ from $(k_{\pi(1)},\dots,k_{\pi(d)})$ where $\pi$ a permutation of $\{1,\dots,d\}$, we obtain a different ${\boldsymbol{z}}$. Therefore, at least $d!$ suitable choices for ${\boldsymbol{z}}$.
The above lemma implies that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}})$, is realized with the prescribed ${\boldsymbol{z}}$ and $n=\#({\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}))$ which is the smallest possible value according to Proposition 4.1. Therefore, in any dimension $d\geq 1$
(4.12) $n_{A}^{*}({\mathcal{B}}_{\boldsymbol{k}})=\#({\mathcal{M}}({\mathcal{B}}_{{\boldsymbol{k}}}))=(2k_{1}+1)\dots(2k_{d}+1).$
5. Lower sets
5.1. Definition and examples
Definition 5.1.
An index set $\Lambda\subset\mathbb{N}_{0}^{d}$ is called lower111also called downward closed if it does satisfy the following property
(5.1) ${\boldsymbol{k}}\in\Lambda\quad\mbox{ and }\quad{\boldsymbol{h}}\leq{\boldsymbol{k}}\quad\implies\quad{\boldsymbol{h}}\in\Lambda.$
The comparison ${\boldsymbol{h}}\leq{\boldsymbol{k}}$ is meant component-wise, $h_{i}\leq k_{i}$ for all $1\leq i\leq d$.
A lower set $\Lambda$ can be equivalently defined by the property
(5.2) ${\boldsymbol{k}}\in\Lambda\quad\text{and}\quad k_{i}\geq 1\quad\implies\quad{\boldsymbol{k}}-{\boldsymbol{e}}_{i}\in\Lambda.$
These particular sets are useful in various applications, as discussed e.g. in [ChkifaLower, Chkifa2016PolynomialAV]. Any lower set contains the null multi-index ${\boldsymbol{0}}$. For a lower set $\Lambda$, we have $\Lambda=\bigcup_{{\boldsymbol{k}}\in\Lambda}{\mathcal{B}}_{\boldsymbol{k}}$. Typical examples of lower sets are the isotropic sets $\left\{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}:k_{j}\leq r\leavevmode\nobreak\ \forall j\right\}$, $\{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}:\sum_{j=1}^{d}k_{j}\leq r\}$, and $\{{\boldsymbol{k}}\in\mathbb{N}_{0}^{d}:\sum_{j=1}^{d}k_{j}^{2}\leq r^{2}\}$, parameterized by $r\geq 1$. These are the intersections of $\mathbb{N}_{0}^{d}$ with balls in $\mathbb{R}^{d}$ of radius $r$, for the $\ell_{\infty}$, $\ell_{1}$ and $\ell_{2}$ norms respectively. Another typical set is the hyperbolic cross $\{{\boldsymbol{k}}:\prod_{j=1}^{d}(1+k_{j})\leq N\}$, which is the union of all lower sets of cardinality less or equal than $N$.
The lower sets we will refer to as of block-type, cross-type or simplex-type, are
(5.3) $\begin{array}[]{c}\displaystyle{\mathcal{B}}_{\boldsymbol{k}}:=\big{\{}{\boldsymbol{h}}\in\mathbb{N}_{0}^{d}:{\boldsymbol{h}}\leq{\boldsymbol{k}}\big{\}},\quad{\mathcal{C}}_{\boldsymbol{k}}:=\big{\{}h_{j}{\boldsymbol{e}}_{j}:0\leq h_{j}\leq k_{j}\big{\}},\quad{\mathcal{S}}_{{\boldsymbol{w}},u}:=\left\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{d}:\sum_{j=1}^{d}w_{j}h_{j}\leq u\right\}.\end{array}$
The first two sets are parameterized by ${\boldsymbol{k}}\in\mathbb{N}_{0}^{d}$ while the latest one by weights $w_{1},\dots,w_{d}>0$ and upper bound $u>0$. These sets are isotropic if ${\boldsymbol{k}}$ (resp. ${\boldsymbol{w}}$) is a multiple of ${\boldsymbol{1}}$. These sets in dimension 2 are illustrated in figure 1.
5.2. Maximal and extremal elements
We say that ${\boldsymbol{k}}$ is a maximal element of $\Lambda$ lower if it is maximal w.r.t. the partial order $\leq$, i.e. ${\boldsymbol{h}}\in\Lambda$ and ${\boldsymbol{k}}\leq{\boldsymbol{h}}$Longrightarrow ${\boldsymbol{h}}={\boldsymbol{k}}$. This is also equivalent to ${\boldsymbol{k}}+{\boldsymbol{e}}_{i}\not\in\Lambda$ for all $i=1,\dots,d$. We say that ${\boldsymbol{k}}$ is an extremal element of $\Lambda$ if ${\boldsymbol{k}}$ cannot be the midpoint of two distinct multi-indices from $\Lambda$, i.e.
(5.4) $2{\boldsymbol{k}}={\boldsymbol{h}}+{\boldsymbol{h}}^{\prime}\quad\text{ and }\quad{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in\Lambda\quad\quad\Longrightarrow\quad\quad{\boldsymbol{h}}={\boldsymbol{h}}^{\prime}={\boldsymbol{k}}.$
The element ${\boldsymbol{0}}$ is clearly extremal. It is also obvious that if ${\boldsymbol{k}}+{\boldsymbol{e}}_{j}\in\Lambda$ for some $j\in\operatorname{supp}({\boldsymbol{k}})$, then ${\boldsymbol{k}}$ is not extremal in $\Lambda$. We denote by ${\mathcal{E}}(\Lambda)$ (resp. ${\mathcal{E}}_{\leq}(\Lambda)$) the set of extremal (resp. maximal) elements of $\Lambda$.
We also introduce notation ${\mathcal{K}}(\Lambda)$ for the elements that are maximal along canonical directions, i.e. ${\boldsymbol{k}}\in{\mathcal{K}}(\Lambda)$ if and only if ${\boldsymbol{k}}=k{\boldsymbol{e}}_{j}\in\Lambda$ for some $k\geq 1$ but ${\boldsymbol{k}}+{\boldsymbol{e}}_{j}\not\in\Lambda$. We note that ${\mathcal{K}}(\Lambda)$ is not necessarily contained in ${\mathcal{E}}_{\leq}(\Lambda)$. Also, a maximal element is not necessarily extremal, however maximal element along canonical directions are.
Any finite lower set has at least one maximal element, and is completely determined by them, i.e.
(5.5) $\Lambda=\bigcup_{{\boldsymbol{k}}\text{ maximal in}\Lambda}{\mathcal{B}}_{\boldsymbol{k}}\;.$
The maximal elements in block/cross/simplex type sets are explicit, i.e.
(5.6) $\begin{array}[]{c}{\mathcal{E}}_{\leq}({\mathcal{B}}_{\boldsymbol{k}})=\big{\{}{\boldsymbol{k}}\big{\}},\quad{\mathcal{E}}_{\leq}({\mathcal{C}}_{\boldsymbol{k}})=\big{\{}k_{j}{\boldsymbol{e}}_{j}:j\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ k_{j}\geq 1\big{\}},\\ {\mathcal{E}}_{\leq}({\mathcal{S}}_{{\boldsymbol{w}},u})=\left\{{\boldsymbol{h}}:u-\min_{i}(w_{i})<\sum_{j=1}^{d}w_{j}h_{j}\leq u\right\},\end{array}$
as well as maximal along canonical directions, i.e.
(5.7) $\begin{array}[]{ccc}{\mathcal{K}}({\mathcal{B}}_{\boldsymbol{k}})={\mathcal{K}}({\mathcal{C}}_{\boldsymbol{k}})={\mathcal{E}}_{\leq}({\mathcal{C}}_{\boldsymbol{k}})&\text{ and }&{\mathcal{K}}({\mathcal{S}}_{{\boldsymbol{w}},u})=\big{\{}\lfloor u/w_{j}\rfloor{\boldsymbol{e}}_{j}:j\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ w_{j}\leq u\big{\}}.\end{array}$
In particular ${\mathcal{E}}_{\leq}({\mathcal{S}}_{{\boldsymbol{1}},k})=\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{d}:\sum_{j=1}^{d}h_{j}=k\}$ and ${\mathcal{K}}({\mathcal{S}}_{{\boldsymbol{1}},k})=\{k{\boldsymbol{e}}_{j}:j=1,\dots,d\}$.
The extremal elements are straightforward for ${\mathcal{B}}_{\boldsymbol{k}}$ and ${\mathcal{C}}_{\boldsymbol{k}}$, i.e.
(5.8) $\begin{array}[]{c}{\mathcal{E}}({\mathcal{B}}_{\boldsymbol{k}})=\big{\{}{\boldsymbol{h}}:h_{j}=0\leavevmode\nobreak\ or\leavevmode\nobreak\ h_{j}=k_{j}\big{\}},\quad{\mathcal{E}}({\mathcal{C}}_{\boldsymbol{k}})=\big{\{}{\boldsymbol{0}}\big{\}}\cup{\mathcal{E}}_{\leq}({\mathcal{C}}_{\boldsymbol{k}}).\end{array}$
For simplex-type sets ${\mathcal{S}}_{{\boldsymbol{w}},u}$, we have ${\mathcal{E}}({\mathcal{S}}_{{\boldsymbol{w}},u})=\{{\boldsymbol{0}}\}\cup{\mathcal{K}}({\mathcal{S}}_{{\boldsymbol{w}},u})\cup{\mathcal{E}}^{(2)}({\mathcal{S}}_{{\boldsymbol{w}},u})$, where ${\mathcal{E}}^{(2)}({\mathcal{S}}_{{\boldsymbol{w}},u})$ consists on extremal elements with at least 2 non-zero components. The characterization of this set is somewhat technical, we simply note the following plain inclusion, if ${\boldsymbol{h}}\in{\mathcal{E}}^{(2)}({\mathcal{S}}_{{\boldsymbol{w}},u})$ then
(5.9) $u-\delta<\sum_{j=1}^{d}w_{j}h_{j}\leq u,\quad\quad\quad\delta=\max_{i}(w_{i})-\min_{i}(w_{i}).$
Indeed, if ${\boldsymbol{h}}\in\mathbb{N}^{d}$ has two non-zero components, say $h_{i},h_{j}\geq 1$, we can write ${\boldsymbol{h}}={\boldsymbol{k}}+{\boldsymbol{k}}^{\prime}$ where ${\boldsymbol{k}}={\boldsymbol{h}}+({\boldsymbol{e}}_{i}-{\boldsymbol{e}}_{j})$ and ${\boldsymbol{k}}^{\prime}={\boldsymbol{h}}+({\boldsymbol{e}}_{j}-{\boldsymbol{e}}_{i})$. We have that ${\boldsymbol{k}}\neq{\boldsymbol{k}}^{\prime}\in\mathbb{N}^{d}$ and $\langle{\boldsymbol{w}},{\boldsymbol{k}}\rangle=\langle{\boldsymbol{w}},{\boldsymbol{h}}\rangle+(w_{i}-w_{j})$, $\langle{\boldsymbol{w}},{\boldsymbol{k}}^{\prime}\rangle=\langle{\boldsymbol{w}},{\boldsymbol{h}}\rangle+(w_{j}-w_{i})$. In particular, if $\langle{\boldsymbol{w}},{\boldsymbol{h}}\rangle\leq u-\delta$, then $\langle{\boldsymbol{w}},{\boldsymbol{k}}\rangle,\langle{\boldsymbol{w}},{\boldsymbol{k}}^{\prime}\rangle\leq u$, implying that ${\boldsymbol{h}}$ is not extremal in ${\mathcal{S}}_{{\boldsymbol{w}},u}$. We note that $\delta$ in (5.9) can be replaced by $\delta_{{\boldsymbol{h}}}=\min_{i\neq j}(|w_{i}-w_{j}|)$ with minimization carried over $i,j\in\operatorname{supp}({\boldsymbol{h}})$.
The same above discussion applied to isotropic simplex, for which $\delta=0$, implies
(5.10) ${\mathcal{E}}({\mathcal{S}}_{{\boldsymbol{1}},k})=\{{\boldsymbol{0}}\}\cup{\mathcal{K}}({\mathcal{S}}_{{\boldsymbol{1}},k}).$
5.3. Sum of lower sets
It is clear that lower structure is preserved by union and intersection. Furthermore, we have the following.
Proposition 5.2.
Let $\Lambda,\Lambda^{\prime}\subset\mathbb{N}_{0}^{d}$ be two lower sets. Then $\Lambda\oplus\Lambda^{\prime}=\{{\boldsymbol{h}}+{\boldsymbol{k}}:{\boldsymbol{h}}\in\Lambda,{\boldsymbol{k}}\in\Lambda^{\prime}\}$ is also lower. Moreover
(5.11) ${\mathcal{M}}(\Lambda\oplus\Lambda^{\prime})={\mathcal{M}}(\Lambda)\oplus{\mathcal{M}}(\Lambda^{\prime}).$
Proof.
First, it is easy to verify that ${\mathcal{B}}_{\boldsymbol{h}}\oplus{\mathcal{B}}_{\boldsymbol{k}}={\mathcal{B}}_{{\boldsymbol{h}}+{\boldsymbol{k}}}$. It is easily verified that $\oplus$ distributes $\cup$, hence $\Lambda\oplus\Lambda^{\prime}=(\cup_{{\boldsymbol{h}}\in\Lambda}{\mathcal{B}}_{\boldsymbol{h}})\oplus(\cup_{{\boldsymbol{k}}\in\Lambda^{\prime}}{\mathcal{B}}_{\boldsymbol{k}})=\cup_{{\boldsymbol{h}}\in\Lambda,{\boldsymbol{k}}\in\Lambda^{\prime}}{\mathcal{B}}_{{\boldsymbol{h}}+{\boldsymbol{k}}}$ is indeed lower. The proof shows incidentally that ${\mathcal{E}}_{\leq}(\Lambda\oplus\Lambda^{\prime})\subset{\mathcal{E}}_{\leq}(\Lambda)\oplus{\mathcal{E}}_{\leq}(\Lambda^{\prime})$. Now, given ${\boldsymbol{h}}+{\boldsymbol{h}}^{\prime}\in\Lambda\oplus\Lambda^{\prime}$ and $s\in\{-1,1\}^{d}$, the multi-index ${\boldsymbol{k}}=s\otimes({\boldsymbol{h}}+{\boldsymbol{h}}^{\prime})=(s_{1}(h_{1}+h_{1}^{\prime}),\dots,s_{d}(h_{d}+h_{d}^{\prime}))$ satisfies ${\boldsymbol{k}}=s\otimes{\boldsymbol{h}}+s\otimes{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda)\oplus{\mathcal{M}}(\Lambda^{\prime})$. As a result ${\mathcal{M}}(\Lambda\oplus\Lambda^{\prime})\subset{\mathcal{M}}(\Lambda)\oplus{\mathcal{M}}(\Lambda^{\prime})$. Inversely, we consider ${\boldsymbol{k}}\in{\mathcal{M}}(\Lambda)$, and ${\boldsymbol{k}}^{\prime}\in{\mathcal{M}}(\Lambda^{\prime})$. We define ${\boldsymbol{h}}$ and ${\boldsymbol{h}}^{\prime}$ in $\mathbb{N}^{d}$ by the following
$\bullet$ $h_{j}=|k_{j}|$ and $h_{j}^{\prime}=|k_{j}^{\prime}|$ if $k_{j}$ and $k_{j}^{\prime}$ have the same signs, otherwise
$\bullet$ $h_{j}=|k_{j}+k_{j}^{\prime}|$ and $h_{j}^{\prime}=0$ if $|k_{j}|\geq|k_{j}^{\prime}|$, or $h_{j}=0$ and $h_{j}^{\prime}=|k_{j}+k_{j}^{\prime}|$ if $|k_{j}|<|k_{j}^{\prime}|$.
By construction, we have $|{\boldsymbol{k}}+{\boldsymbol{k}}^{\prime}|={\boldsymbol{h}}+{\boldsymbol{h}}^{\prime}$ with ${\boldsymbol{h}}\leq|{\boldsymbol{k}}|$ and ${\boldsymbol{h}}^{\prime}\leq|{\boldsymbol{k}}^{\prime}|$. Since $\Lambda$ and $\Lambda^{\prime}$ are lower, then ${\boldsymbol{h}}$ belongs to $\Lambda$ and ${\boldsymbol{h}}^{\prime}$ belongs to $\in\Lambda^{\prime}$, implying that $|{\boldsymbol{k}}+{\boldsymbol{k}}^{\prime}|$ belongs to $\Lambda\oplus\Lambda$. Therefore ${\boldsymbol{k}}+{\boldsymbol{k}}^{\prime}$ belongs to ${\mathcal{M}}(\Lambda\oplus\Lambda)$. The proof is complete. ∎
As mentioned above, the sum $\oplus$ of block-type sets is plain, i.e. ${\mathcal{B}}_{\boldsymbol{h}}\oplus{\mathcal{B}}_{\boldsymbol{k}}={\mathcal{B}}_{{\boldsymbol{h}}+{\boldsymbol{k}}}$. For cross-type sets, we have ${\mathcal{C}}_{\boldsymbol{h}}\oplus{\mathcal{C}}_{\boldsymbol{k}}=\cup_{1\leq i,j\leq d}\leavevmode\nobreak\ {\mathcal{B}}_{h_{i}{\boldsymbol{e}}_{i}+k_{j}{\boldsymbol{e}}_{j}}$. As for simplex-type sets,
(5.12) ${\mathcal{S}}_{{\boldsymbol{w}},u_{1}+u_{2}-\max_{i}(w_{i})}\subset{\mathcal{S}}_{{\boldsymbol{w}},u_{1}}\oplus{\mathcal{S}}_{{\boldsymbol{w}},u_{2}}\subset{\mathcal{S}}_{{\boldsymbol{w}},u_{1}+u_{2}}.$
[…]
Next, we establish an inequality for lower sets in terms of the size of their mirrored and sum sets.
Proposition 5.3.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ a finite lower set. There holds
(5.13) #$\#({\mathcal{M}}(\Lambda))=\sum_{{\boldsymbol{k}}\in\Lambda}2^{|{\boldsymbol{k}}|_{0}}\leq\#(\Lambda\oplus\Lambda).$
Equality holds if and only if $\Lambda={\mathcal{B}}_{{\boldsymbol{k}}}$ for some ${\boldsymbol{k}}\in\mathbb{N}_{0}^{d}$.
Proof.
For a multi-index ${\boldsymbol{k}}\in\mathbb{N}_{0}^{d}$, we let ${\rm supp}({\boldsymbol{k}})$ be the support of ${\boldsymbol{k}}$, which is the subset of $\{1,\dots,d\}$ consisting of $j$ s.t. $k_{j}\neq 0$. We define ${\mathcal{D}}_{{\boldsymbol{k}}}$ to be the set of $\varepsilon\in\{0,1\}^{d}$ satisfying ${\rm supp}(\varepsilon)\subset{\rm supp}({\boldsymbol{k}})$. We then consider a set-valued map over $\mathbb{N}_{0}^{d}$ by
$F:{\boldsymbol{k}}\longmapsto\{2{\boldsymbol{k}}-\varepsilon\}_{\varepsilon\in{\mathcal{D}}_{{\boldsymbol{k}}}}.$
This map satisfies $F({\boldsymbol{k}})\cap F({\boldsymbol{k}}^{\prime})=\emptyset$ if ${\boldsymbol{k}}\neq{\boldsymbol{k}}^{\prime}$. First, it is easily verified $F({\boldsymbol{0}})=\{{\boldsymbol{0}}\}$ and ${\boldsymbol{0}}$ does not belong to any other $F({\boldsymbol{k}})$. If we are given ${\boldsymbol{k}},{\boldsymbol{k}}^{\prime}\in\mathbb{N}_{0}^{d}-\{{\boldsymbol{0}}\}$ s.t. $F({\boldsymbol{k}})\cap F({\boldsymbol{k}}^{\prime})\neq\emptyset$, then for some $\varepsilon\in{\mathcal{D}}_{{\boldsymbol{k}}}$ and $\varepsilon^{\prime}\in{\mathcal{D}}_{{\boldsymbol{k}}^{\prime}}$ we must have $2{\boldsymbol{k}}-\varepsilon=2{\boldsymbol{k}}^{\prime}-\varepsilon^{\prime}$. But this implies $2({\boldsymbol{k}}-{\boldsymbol{k}}^{\prime})=\varepsilon-\varepsilon^{\prime}$ belongs to $\{-1,0,1\}^{d}$, which is possible only if ${\boldsymbol{k}}-{\boldsymbol{k}}^{\prime}={\boldsymbol{0}}$. Now, given $\Lambda$ lower, we have for any ${\boldsymbol{k}}\in\Lambda$ and any $\varepsilon\in{\mathcal{D}}_{{\boldsymbol{k}}}$ that $2{\boldsymbol{k}}-\varepsilon={\boldsymbol{k}}+({\boldsymbol{k}}-\varepsilon)$ belongs to $\Lambda\oplus\Lambda$. Therefore $\cup_{{\boldsymbol{k}}\in\Lambda}F({\boldsymbol{k}})\subset\Lambda\oplus\Lambda$. This combined with empty intersection property implies
#$\#(\Lambda\oplus\Lambda)\geq\sum_{{\boldsymbol{k}}\in\Lambda}\#(\{2{\boldsymbol{k}}-\varepsilon\}_{\varepsilon\in{\mathcal{D}}_{{\boldsymbol{k}}}})=\sum_{{\boldsymbol{k}}\in\Lambda}\#({\mathcal{D}}_{{\boldsymbol{k}}})=\sum_{{\boldsymbol{k}}\in\Lambda}2^{|{\boldsymbol{k}}|_{0}}=\#({\mathcal{M}}(\Lambda)).$
The first part of the proposition is complete. We note that $F(\Lambda):=\cup_{{\boldsymbol{k}}\in\Lambda}F({\boldsymbol{k}})$ has an explicit formula for $\Lambda$ a lower set, i.e. $F(\Lambda)=\cup_{{\boldsymbol{k}}\in\Lambda}{\mathcal{B}}_{2{\boldsymbol{k}}}=\cup_{{\boldsymbol{k}}\in\Lambda\text{ maximal }}{\mathcal{B}}_{2{\boldsymbol{k}}}$. On the other hand, we have $\Lambda\oplus\Lambda=\cup_{{\boldsymbol{h}},{\boldsymbol{k}}\in\Lambda}{\mathcal{B}}_{{\boldsymbol{h}}+{\boldsymbol{k}}}=\cup_{{\boldsymbol{h}},{\boldsymbol{k}}\in\Lambda\text{ maximal }}{\mathcal{B}}_{{\boldsymbol{h}}+{\boldsymbol{k}}}$. Clearly, equality holds for block-type sets, with $F({\mathcal{B}}_{\boldsymbol{h}})={\mathcal{B}}_{2{\boldsymbol{h}}}={\mathcal{B}}_{{\boldsymbol{h}}}\oplus{\mathcal{B}}_{{\boldsymbol{h}}}$. In general, it seems unlikely that for all pairs $({\boldsymbol{i}},{\boldsymbol{j}})$ of maximal elements of $\Lambda$, one has ${\boldsymbol{i}}+{\boldsymbol{j}}\subset F(\Lambda)$. We give a proof to this observation.
We let $\Lambda$ be lower not of block-type and assume $\Lambda\oplus\Lambda=F(\Lambda)$. Let ${\boldsymbol{i}}^{(1)},\dots,{\boldsymbol{i}}^{(m)}$ be its pairwise different maximal elements ordered in lexicographical order, ${\boldsymbol{h}}\leq_{L}{\boldsymbol{k}}$ if and only if $h_{1}<k_{1}$ or ($h_{1}=k_{1}$ and $\widehat{\boldsymbol{h}}\leq_{L}\widehat{\boldsymbol{k}}$) where $\widehat{\boldsymbol{x}}=(x_{2},\dots,x_{d})$. By the assumption $\Lambda\oplus\Lambda=F(\Lambda)$, we must have ${\boldsymbol{i}}^{(m-1)}+{\boldsymbol{i}}^{(m)}\leq 2{\boldsymbol{i}}^{(n)}$ for some $n\in\{1,\dots,m\}$. We cannot have $n=m-1$ or $n=m$ since since then either ${\boldsymbol{i}}^{(m)}$ or ${\boldsymbol{i}}^{(m-1)}$ cannot be maximal elements which is in contradiction to the assumption. If $m=2$, we reach already a contradiction. We assume that $m\geq 3$ and use notation $\nu={\boldsymbol{i}}^{(m-1)}$, $\mu={\boldsymbol{i}}^{(m)}$, $\lambda={\boldsymbol{i}}^{(n)}$, hence $\nu+\mu\leq 2\lambda$. Using the lexicographical order assumption, at position $1$ we must have $\lambda_{1}\leq\nu_{1}$ and $\lambda_{1}\leq\mu_{1}$, hence $2\lambda_{1}\leq\nu_{1}+\mu_{1}$. If one of the first inequalities is strict, we get $2\lambda_{1}<\nu_{1}+\mu_{1}$ contradicting $\nu+\mu\leq 2\lambda$. Re-iterating this argument shows that we must have $\lambda_{j}=\nu_{j}$ and $\lambda_{j}=\mu_{j}$ for all $j$. We again reach a contradiction. The proof is complete. ∎
With little extra effort, we identify a map that allows to embed ${\mathcal{M}}(\Lambda)$ into $\Lambda\oplus\Lambda$ for any lower set $\Lambda$. In view of the application $F$ defined in the proof, the following map satisfies the identity
(5.14) ${\boldsymbol{k}}\in\mathbb{Z}^{d}\mapsto 2|{\boldsymbol{k}}|-{\bf 1}_{{\boldsymbol{k}}<{\boldsymbol{0}}}\in\mathbb{N}_{0}^{d},$
where $|{\boldsymbol{k}}|$ is component-wise absolute value of ${\boldsymbol{k}}$ and ${\bf 1}_{{\boldsymbol{k}}<{\boldsymbol{0}}}\in\{0,1\}^{d}$ having ones in positions $j$ s.t. $k_{j}<0$.
6. Equivalence between reconstruction plans for lower sets
In this section, we establish some equivalence properties between admissibility requirements (3.16), (3.21) (3.22) and (3.23) when $\Lambda$ is lower and more particularly when $\Lambda$ is of simplex-type.
6.1. Equivalences for general lower sets
We start with the following result.
Proposition 6.1.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be a lower set, $n\in\mathbb{N}$, and ${\boldsymbol{z}}\in\mathbb{Z}^{d}$. The following assertions are equivalent:
  • $(n,{\boldsymbol{z}})$ is admissible for plan 0 for $\Lambda\oplus\Lambda$,
  • $(n,{\boldsymbol{z}})$ is admissible for plan A for $\Lambda$.
Indeed, we have by proposition 5.2 that ${\mathcal{M}}(\Lambda\oplus\Lambda)={\mathcal{M}}(\Lambda)\oplus{\mathcal{M}}(\Lambda)={\mathcal{M}}(\Lambda)\ominus{\mathcal{M}}(\Lambda)$. In particular, ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0$ for all ${\boldsymbol{h}}\in{\mathcal{M}}(\Lambda\oplus\Lambda)\setminus\{{\boldsymbol{0}}\}$ is equivalent to ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}},{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda)$ s.t. ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}$.
This next result is also straightforward.
Proposition 6.2.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be a lower set, $n\in\mathbb{N}$ and ${\boldsymbol{z}}\in\mathbb{Z}^{d}$. Then,
(6.1) ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda\quad\quad\implies\quad\quad{\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)\setminus\{0\}.$
An equivalence holds if $\Lambda={\mathcal{B}}_{\boldsymbol{k}}$ for some ${\boldsymbol{k}}\in\mathbb{N}_{0}^{d}$.
Proof.
Given any ${\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)\setminus\{{\boldsymbol{0}}\}$, we can decompose it as ${\boldsymbol{h}}={\boldsymbol{h}}^{+}-{\boldsymbol{h}}^{-}$ where ${\boldsymbol{h}}^{+}=\max({\boldsymbol{h}},{\boldsymbol{0}})$ and ${\boldsymbol{h}}^{-}=\max(-{\boldsymbol{h}},{\boldsymbol{0}})$ both in $\mathbb{N}^{d}_{0}$. Since ${\boldsymbol{h}}^{+},{\boldsymbol{h}}^{-}\leq|{\boldsymbol{h}}|$, $|{\boldsymbol{h}}|\in\Lambda$, and $\Lambda$ is lower, then ${\boldsymbol{h}}^{+},{\boldsymbol{h}}^{-}\in\Lambda$. Also by construction ${\boldsymbol{h}}^{+}\neq{\boldsymbol{h}}^{-}$, and ${\boldsymbol{h}}^{-}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{+}\cdot{\boldsymbol{z}}$ implies ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0$. If $\Lambda={\mathcal{B}}_{k}$, we use that ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in\Lambda$ is equivalent to ${\boldsymbol{0}}\neq({\boldsymbol{h}}-{\boldsymbol{h}}^{\prime})\in{\mathcal{M}}(\Lambda)$. ∎
The following results are more involved, however rely on the same techniques as above. First, we address the equivalence between plans B and C, i.e. requirements (3.22) and (3.23).
Proposition 6.3.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be a lower set, $n\in\mathbb{N}$, and ${\boldsymbol{z}}\in\mathbb{Z}^{d}$. The following assertions are equivalent:
  • $(n,{\boldsymbol{z}})$ is admissible for plan B,
  • $(n,{\boldsymbol{z}})$ is admissible for plan C and $02{\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ \not\equiv_{n}$ 0 for all ${\boldsymbol{h}}\in{\mathcal{E}}(\Lambda)\setminus\{0\}$.
Moreover, if $n$ is odd, the last assumption is not needed.
Proof.
It is immediate that plan B admissibility does imply plan C admissibility and does imply that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ \not\equiv_{n}-{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\in\Lambda\setminus\{{\boldsymbol{0}}\}$, see Remark 3.5. As for the other implication, we use proof by contradiction. We assume that $(n,{\boldsymbol{z}})$ is admissible for plan C and satisfies the identity but $(n,{\boldsymbol{z}})$ is not admissible for plan B. In view of Remark 3.5, there must exist ${\boldsymbol{h}}\in\Lambda\setminus\{{\boldsymbol{0}}\}$ and $\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}^{*}$ such that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}$. We have that $\sigma({\boldsymbol{h}})={\boldsymbol{h}}^{+}-{\boldsymbol{h}}^{-}$ and ${\boldsymbol{h}}={\boldsymbol{h}}^{+}+{\boldsymbol{h}}^{-}$ where ${\boldsymbol{h}}^{+}=\max(\sigma({\boldsymbol{h}}),{\boldsymbol{0}})$ and ${\boldsymbol{h}}^{-}=\max(-\sigma({\boldsymbol{h}}),{\boldsymbol{0}})$ both in $\mathbb{N}^{d}_{0}$. Let us note that ${\boldsymbol{h}}^{-}\neq{\boldsymbol{0}}$, since ${\boldsymbol{h}}^{-}={\boldsymbol{0}}$ is only possible if $\sigma=id$ which is excluded. The equality ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}$ is reformulated
$2{\boldsymbol{h}}^{-}\cdot{\boldsymbol{z}}\equiv_{n}0\equiv_{n}{\boldsymbol{0}}\cdot{\boldsymbol{z}}.$
In particular, if $2{\boldsymbol{h}}^{-}\in\Lambda$ we reach a contradiction. In general, since $\Lambda$ is lower then ${\boldsymbol{h}}^{-}\in\Lambda$. If ${\boldsymbol{h}}^{-}\not\in{\mathcal{E}}(\Lambda)$, then $2{\boldsymbol{h}}^{-}={\boldsymbol{k}}+{\boldsymbol{k}}^{\prime}$ for some ${\boldsymbol{k}}\neq{\boldsymbol{k}}^{\prime}\in\Lambda$. Then $02{\boldsymbol{h}}^{-}\cdot{\boldsymbol{z}}\equiv_{n}$ 0 implies ${\boldsymbol{k}}\cdot{\boldsymbol{z}}=-{\boldsymbol{k}}^{\prime}\cdot{\boldsymbol{z}}$, which contradicts plan C admissibility. The proof of the equivalence is complete. If $n$ is odd, then $02{\boldsymbol{h}}^{-}\cdot{\boldsymbol{z}}\not\equiv_{n}$ 0 is equivalent to ${\boldsymbol{h}}^{-}\cdot{\boldsymbol{z}}\not\equiv_{n}0$, but this is already incorporated in plan C admissibility since ${\boldsymbol{0}}\neq{\boldsymbol{h}}^{-}\in\Lambda$. The proof is complete.
In general, the requirement $02{\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ \not\equiv_{n}$ 0 for all ${\boldsymbol{h}}\in{\mathcal{E}}(\Lambda)\setminus\{0\}$ seems quite involved since it entails identifying the set of extremal elements ${\mathcal{E}}(\Lambda)$. One can therefore always favor searching for $n$ odd. Since $n^{*}_{C}(\Lambda)\leq n^{*}_{B}(\Lambda)$ and $n^{*}_{C}(\Lambda,{\boldsymbol{z}})\leq n^{*}_{B}(\Lambda,{\boldsymbol{z}})$ for any set $\Lambda\subset\mathbb{N}_{0}^{d}$ and for any ${\boldsymbol{z}}$, the following holds when we choose $\Lambda$ to be a lower set,
  • $n^{*}_{C}(\Lambda,{\boldsymbol{z}})$ is odd $\implies$ $n^{*}_{B}(\Lambda,{\boldsymbol{z}})=n^{*}_{C}(\Lambda,{\boldsymbol{z}})$,
  • $n^{*}_{C}(\Lambda)$ is odd $\implies$ $n^{*}_{B}(\Lambda)=n^{*}_{C}(\Lambda)$.
Remark 6.4.
In the general case, we cannot ensure $n^{*}_{C}(\Lambda)=n^{*}_{B}(\Lambda)$, as per the case $\Lambda=\{0,1\}$, which is a lower set in dimension $1$. For plan C, we need $z\not\equiv_{n}0$, hence $(n=2,{\boldsymbol{z}}=1)$ is admissible and $n^{*}_{C}(\Lambda)=2$. For plan B, we need $z\not\equiv_{n}0$ and $z\not\equiv_{n}-z$ which implies that $(n=3,{\boldsymbol{z}}=1)$ is admissible and $n^{*}_{B}(\Lambda)=3$.
Remark 6.5.
Given any $\Lambda\subset\mathbb{N}_{0}^{d}$ lower and $n\in\mathbb{N}$ odd, having ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ for some ${\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)$ both belonging to the same $\{\sigma({\boldsymbol{k}}),\sigma\in{\mathcal{S}}_{\boldsymbol{k}}\}$ is equivalent to $\widetilde{\boldsymbol{h}}\cdot{\boldsymbol{z}}\equiv_{n}0$ where $\widetilde{\boldsymbol{h}}\neq{\boldsymbol{0}}$ is defined by $\widetilde{h}_{j}=k_{j}$ if $h_{j}^{\prime}=-h_{j}$ and $\widetilde{h}_{j}=0$ otherwise. Since $\widetilde{\boldsymbol{h}}\in\Lambda$, this is ruled out if plan C admissibility is assumed. In light of this observation, the table of Remark 3.5 can be simplified for $\Lambda$ lower and $n\in\mathbb{N}$ odd to
$\begin{array}[]{|l l|c|c|c|}\hline\cr\text{Requirement}&&\text{C}&\text{ B}&\text{A}\\ \hline\cr{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}&\quad{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\leavevmode\nobreak\ \sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}&\checkmark&\checkmark&\checkmark\\ {\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}&\quad{\boldsymbol{h}}^{\prime},{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda),|{\boldsymbol{h}}^{\prime}|\neq|{\boldsymbol{h}}|&&&\checkmark\\ \hline\cr\end{array}$
6.2. Equivalence for the isotropic simplex-type sets
We now specialize to simplex-type sets. For these sets, minimal conditions are required in order to have equivalence between plans. First, for ${\mathcal{S}}_{{\boldsymbol{1}},k}$ by simply combining Proposition 6.3 with the fact that ${\mathcal{E}}({\mathcal{S}}_{{\boldsymbol{1}},k})\setminus\{{\boldsymbol{0}}\}={\mathcal{K}}({\mathcal{S}}_{{\boldsymbol{1}},k})=\{k{\boldsymbol{e}}_{j}:j=1,\dots,d\}$, see (5.10), we have the equivalence between these two assertions:
  • $(n,{\boldsymbol{z}})$ is admissible for plan B,
  • $(n,{\boldsymbol{z}})$ is admissible for plan C and $02kz_{j}\not\equiv_{n}$ 0 for all $j=1,\dots,d$,
with the last assumption not needed if $n$ is odd. A stronger result actually holds.
Proposition 6.6.
Let $k,n\in\mathbb{N}$, and ${\boldsymbol{z}}\in\mathbb{Z}^{d}$. For $\Lambda={\mathcal{S}}_{{\boldsymbol{1}},k}$, the following assertions are equivalent:
  • $(n,{\boldsymbol{z}})$ is admissible for plan A
  • $(n,{\boldsymbol{z}})$ is admissible for plan B
  • $(n,{\boldsymbol{z}})$ is admissible for plan C and $02kz_{j}\not\equiv_{n}$ 0 for all $j=1,\dots,d$.
Moreover, if $n$ is odd, the last assumption is not needed.
Proof.
The forward implications are immediate. We only need to prove the reciprocate ones. We assume $(n,{\boldsymbol{z}})$ is admissible for plan C and $02kz_{j}\not\equiv_{n}$ 0 for all $j=1,\dots,d$, but not admissible for plan A. Let ${\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in{\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k})$ be s.t. ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}$. We decompose ${\boldsymbol{h}}={\boldsymbol{h}}^{+}-{\boldsymbol{h}}^{-}$ and ${\boldsymbol{h}}^{\prime}={\boldsymbol{h}}^{\prime+}-{\boldsymbol{h}}^{\prime-}$, leading to ${\boldsymbol{k}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{k}}^{\prime}\cdot{\boldsymbol{z}}$ where ${\boldsymbol{k}}={\boldsymbol{h}}^{+}+{\boldsymbol{h}}^{\prime-}$ and ${\boldsymbol{k}}^{\prime}={\boldsymbol{h}}^{\prime+}+{\boldsymbol{h}}^{-}$, which both belong to $\mathbb{N}_{0}^{d}$ and satisfy ${\boldsymbol{k}}\neq{\boldsymbol{k}}^{\prime}$. Also, up to subtract $\min({\boldsymbol{k}},{\boldsymbol{k}}^{\prime})$ from ${\boldsymbol{k}}$ and ${\boldsymbol{k}}^{\prime}$, and swap them if necessary, we may assume that $\operatorname{supp}({\boldsymbol{k}})\cap\operatorname{supp}({\boldsymbol{k}}^{\prime})=\emptyset$ and $\|{\boldsymbol{k}}\|_{\ell_{1}}\leq\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}$. We have
$2\|{\boldsymbol{k}}\|_{\ell_{1}}\leq\|{\boldsymbol{k}}\|_{\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}=\|{\boldsymbol{h}}\|_{\ell_{1}}+\|{\boldsymbol{h}}^{\prime}\|_{\ell_{1}}\leq 2k,$
hence ${\boldsymbol{k}}\in{\mathcal{S}}_{{\boldsymbol{1}},k}$. If ${\boldsymbol{k}}^{\prime}\in{\mathcal{S}}_{{\boldsymbol{1}},k}$, we reach a contradiction with plan C admissibility. We consider the case ${\boldsymbol{k}}^{\prime}\not\in{\mathcal{S}}_{{\boldsymbol{1}},k}$, namely $\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}>k$. Without loss of generality, we assume that $\operatorname{supp}({\boldsymbol{k}}^{\prime})=\{1,\dots,J\}$. We then let ${\boldsymbol{y}}={\boldsymbol{k}}$ and ${\boldsymbol{y}}^{\prime}={\boldsymbol{k}}^{\prime}$, loop over $j=1,\dots,J$ and within every loop, while $\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}>k$ do ${\boldsymbol{y}}\leftarrow{\boldsymbol{y}}-{\boldsymbol{e}}_{j}$ and ${\boldsymbol{y}}^{\prime}\leftarrow{\boldsymbol{y}}^{\prime}-{\boldsymbol{e}}_{j}$. This procedure produces ${\boldsymbol{y}}\in\mathbb{Z}^{d}$ and ${\boldsymbol{y}}^{\prime}\in\mathbb{N}_{0}^{d}$ satisfying $\|{\boldsymbol{y}}\|_{\ell_{1}}+\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}=\|{\boldsymbol{k}}\|_{\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}$ and $\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}=k$. In particular ${\boldsymbol{y}}\in{\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{1}},k})$ and ${\boldsymbol{y}}^{\prime}\in{\mathcal{S}}_{{\boldsymbol{1}},k}$. Since ${\boldsymbol{y}}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{y}}^{\prime}\cdot{\boldsymbol{z}}$, then if $|{\boldsymbol{y}}|\neq{\boldsymbol{y}}^{\prime}$ we reach a contradiction with plan C admissibility. We observe that $|{\boldsymbol{y}}|\neq{\boldsymbol{y}}^{\prime}$ will always hold unless $J=1$, ${\boldsymbol{k}}={\boldsymbol{0}}$ and ${\boldsymbol{k}}^{\prime}=2ke_{1}$.
In this case ${\boldsymbol{k}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{k}}\cdot{\boldsymbol{z}}$ is reformulated as $02kz_{1}\equiv_{n}$ 0,
hence a contradiction with the assumption on $k$ and ${\boldsymbol{z}}$.
When $n$ is odd, the assumption $02kz_{1}\not\equiv_{n}0,\dots,2kz_{d}\not\equiv_{n}$ 0 reduces to $kz_{1}\not\equiv_{n}$,…, $kz_{d}\not\equiv_{n}0$ which is incorporated in plan C admissibility. The proof is complete. ∎
The arguments used in the proof can be used in order to simplify the requirements. In the first case of the proof, i.e. $\|{\boldsymbol{k}}\|_{\ell_{1}}\leq\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}\leq k$, we can define ${\boldsymbol{y}}={\boldsymbol{k}}+\delta e_{1}$ and ${\boldsymbol{y}}^{\prime}={\boldsymbol{k}}^{\prime}+\delta e_{1}$ with $\delta=(k-\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}})$, so that $\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}=k$. This identity is also satisfied with ${\boldsymbol{y}}^{\prime}$ produced by the procedure in the second case. In view of this, we can verify that for $\Lambda={\mathcal{S}}_{{\boldsymbol{1}},k}$, $(n,{\boldsymbol{z}})$ is admissible for plan C if and only if
(6.2) ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}},\quad\mbox{for}\quad\begin{array}[]{c}{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\quad\quad\|{\boldsymbol{h}}^{\prime}\|_{\ell^{1}}=k.\end{array}$
Likewise, we can carry on the procedure updating ${\boldsymbol{y}}$ and ${\boldsymbol{y}}^{\prime}$, while $\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}>\|{\boldsymbol{y}}\|_{\ell_{1}}+1$, instead of stopping when $\|{\boldsymbol{y}}^{\prime}\|_{\ell_{1}}=k$, and use this as well if $\|{\boldsymbol{k}}\|_{\ell_{1}}\leq\|{\boldsymbol{k}}^{\prime}\|_{\ell_{1}}\leq k$. We establish that for $\Lambda={\mathcal{S}}_{{\boldsymbol{1}},k}$, $(n,{\boldsymbol{z}})$ is admissible for plan C if and only if
(6.3) ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}},\quad\mbox{for}\quad\begin{array}[]{c}{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\quad\quad\|{\boldsymbol{h}}^{\prime}\|_{\ell^{1}}-\|{\boldsymbol{h}}\|_{\ell^{1}}\in\{0,1\}.\end{array}$
6.3. Equivalence for the anisotropic simplex-type sets
The techniques we have used in the proof of Proposition 6.6 can be re-examined with general simplex-type sets. For $w_{1},\dots,w_{d}>0$, one has that $\|{\boldsymbol{x}}\|_{{\boldsymbol{w}},\ell_{1}}:=\sum_{j=1}^{d}w_{j}|x_{j}|$ defines a norm over $\mathbb{R}^{d}$, and that ${\mathcal{S}}_{{\boldsymbol{w}},u}=\{{\boldsymbol{k}}\in\mathbb{N}^{d}:\|{\boldsymbol{k}}\|_{{\boldsymbol{w}},\ell_{1}}\leq u\}$, ${\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{w}},u})=\{{\boldsymbol{h}}\in\mathbb{Z}^{d}:\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}}\leq u\}$. The proof is unchanged if $\|\cdot\|_{\ell_{1}}$ is replaced by $\|\cdot\|_{{\boldsymbol{w}},\ell_{1}}$ and $k$ is replaced by $u$, up until we consider ${\boldsymbol{k}}^{\prime}\not\in{\mathcal{S}}_{{\boldsymbol{w}},u}$. Then we distinguish two cases
  • $\|{\boldsymbol{k}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq 2u-\alpha$, and
  • $\|{\boldsymbol{k}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}>2u-\alpha$,
where $\alpha=\max_{i\in\operatorname{supp}({\boldsymbol{k}}^{\prime})}(w_{i})$. In the first case, the same procedure as the proof yields ${\boldsymbol{y}}$ and ${\boldsymbol{0}}\leq{\boldsymbol{y}}^{\prime}\leq{\boldsymbol{k}}^{\prime}$ satisfying $u-\alpha<\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$, and $\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}=\|{\boldsymbol{k}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq 2u-\alpha$, hence $\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$. Also $|{\boldsymbol{y}}|\neq{\boldsymbol{y}}^{\prime}$ unless $J=1$, ${\boldsymbol{k}}={\boldsymbol{0}}$ and ${\boldsymbol{k}}^{\prime}=2\lfloor u/w_{1}\rfloor e_{1}$ assuming $w_{1}\leq u$. The second case originates in the proof from ${\boldsymbol{h}}$ and ${\boldsymbol{h}}^{\prime}$ s.t. $\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}>2u-\alpha$.
We introduce notation $\alpha_{{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}}:=\max(w_{i})$ with minimization carried over $i\in\operatorname{supp}({\boldsymbol{h}})\cup\operatorname{supp}({\boldsymbol{h}}^{\prime})$. In view of the above discussion, we identify the equivalence presented in the following proposition.
Proposition 6.7.
Let $w_{1},\dots,w_{d},u>0$, and $n\in\mathbb{N}$, ${\boldsymbol{z}}\in\mathbb{Z}^{d}$. For $\Lambda={\mathcal{S}}_{{\boldsymbol{w}},u}$, the following are equivalent:
  • $(n,{\boldsymbol{z}})$ is admissible for plan A
  • $(n,{\boldsymbol{z}})$ is admissible for plan C, $02\lfloor u/w_{j}\rfloor z_{j}\not\equiv_{n}$ 0 for $j$ s.t. $u\geq w_{j}$, and
    (6.4) ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}\quad\mbox{for}\quad\begin{array}[]{c}{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)\quad\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}>2u-\alpha_{{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}}.\end{array}$
If $n$ is odd, the second assumption is not needed.
Characterizations such as (6.2) and (6.3) can be derived for general simplex-type sets. We observe that if ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}}$ for some ${\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in{\mathcal{M}}({\mathcal{S}}_{{\boldsymbol{w}},u})$, we may assume that ${\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{E}}_{\leq}({\mathcal{S}}_{{\boldsymbol{w}},u}))$. Indeed, up to multiply by $-1$ and swap if necessary, we may assume $\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}}\leq\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}$ and $h_{i}^{\prime}\geq 0$ where $i$ is s.t. $w_{i}$ is minimal among the $w_{j}$. Then by considering ${\boldsymbol{y}}={\boldsymbol{h}}+\delta{\boldsymbol{e}}_{i}$ and ${\boldsymbol{y}}^{\prime}={\boldsymbol{h}}^{\prime}+\delta{\boldsymbol{e}}_{i}$ where $\delta$ is the largest integer s.t. $|{\boldsymbol{y}}^{\prime}|\in{\mathcal{S}}_{{\boldsymbol{w}},u}$, we obtain $|{\boldsymbol{y}}^{\prime}|$ is maximal and $\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}\leq\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$. We therefore have for $\Lambda={\mathcal{S}}_{{\boldsymbol{w}},u}$ that $(n,{\boldsymbol{z}})$ is admissible for plan A if and only if
(6.5) ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}\cdot{\boldsymbol{z}},\quad\mbox{ for }\quad{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda),\;\;{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}({\mathcal{E}}_{\leq}(\Lambda)),\;\;{\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}.$
In procedure we discussed above, if we carry on updating ${\boldsymbol{y}}$ and ${\boldsymbol{y}}^{\prime}$ while $\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}-\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}>\alpha$, instead of stopping when $\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$, we obtain that $|\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}-\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}|\leq\alpha$, also implying $\|{\boldsymbol{y}}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$ and $\|{\boldsymbol{y}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$. The other case $\|{\boldsymbol{k}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{k}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}>2u-\alpha$, which results of $\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}}+\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}>2u-\alpha$, implies that ${\boldsymbol{h}}$ and ${\boldsymbol{h}}^{\prime}$ are necessarily s.t. $u-\alpha<\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell_{1}},\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell_{1}}\leq u$. In particular, we identify that for sets $\Lambda={\mathcal{S}}_{{\boldsymbol{w}},u}$, $(n,{\boldsymbol{z}})$ is admissible for plan C if and only if
(6.6) ${\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\not\equiv_{n}\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}},\quad\mbox{ for }\quad{\boldsymbol{h}}^{\prime}\neq{\boldsymbol{h}}\in\Lambda,\;\;\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}},\;\;\left|\|{\boldsymbol{h}}^{\prime}\|_{{\boldsymbol{w}},\ell^{1}}-\|{\boldsymbol{h}}\|_{{\boldsymbol{w}},\ell^{1}}\right|<\alpha_{{\boldsymbol{h}},{\boldsymbol{h}}^{\prime}}.$
7. Specialized CBC construction
In this section, we introduce efficient numerical algorithms for determining moderate values of $n$ and the generator vector ${\boldsymbol{z}}$ for any given lower set $\Lambda$. Our algorithm offers substantial improvements over current methods in both computation time and memory efficiency, making it particularly effective for high-dimensional cases.
7.1. Exhaustive search
We let $\Lambda\subset\mathbb{N}_{0}^{d}$ be an arbitrary set of finite cardinality and $(n,{\boldsymbol{z}})\in\mathbb{N}\times\mathbb{Z}^{d}$. We recall that $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{0}(\Lambda)$ iff (3.16) holds. On the other hand, $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{A}(\Lambda)$ (resp. $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{B}(\Lambda)$, $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{C}(\Lambda)$) iff (3.21) holds (resp. (3.22) holds, (3.23) holds).
The goal is to find a couple $(n,{\boldsymbol{z}})$ consisting in a small integer $n$ and ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ s.t. $({\boldsymbol{z}},n)\in{\mathcal{P}}_{\square}(\Lambda)$ with $\square$ is a placeholder for 0, A, B or C. Exhaustive search entails incrementally testing all possible combinations of $n$ and ${\boldsymbol{z}}$ ($\in\{0,\dots,n-1\}^{d}$) until the admissibility condition is met. A unified implementation of the search is Algorithm 1.
   Inputs: $\Lambda\subset\mathbb{N}^{d}$, and $n_{\rm lower}$, $n_{\rm upper}$ theoretical lower and upper bound on $n$
  $n=n_{\rm lower}$
  while $n\leq n_{\rm upper}$ do
     for ${\boldsymbol{z}}\in\{0,\dots,n-1\}^{d}$ do
        if  $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{\square}(\Lambda)$ then
           return  $n,{\boldsymbol{z}}$
        end if
     end for
     Increment $n$ by 1
  end while
Algorithm 1 Exhaustive search for optimal $n$ and an associated generator ${\boldsymbol{z}}$
Let us comment on the algorithm. In light of equations (3.24) and (4.4), whatever the geometry of $\Lambda$, one has closed formulas for $n_{\rm lower}$ and $n_{\rm upper}$. Checking whether $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{\square}(\Lambda)$ for $\Lambda$ entails iterating through elements ${\boldsymbol{h}}$ in ${\mathcal{M}}(\Lambda)$ and making sure that the inner products ${\boldsymbol{h}}\cdot{\boldsymbol{z}}$ satisfy a prescribed condition, as summarized in Remark 3.5. More precisely
  • integral exactness: verify that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}0$ if ${\boldsymbol{h}}\neq{\boldsymbol{0}}$,
  • reconstruction with plan A: verify that $({\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ are pairwise distinct,
  • reconstruction with plans B/C: first list all $({\boldsymbol{h}}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ for ${\boldsymbol{h}}$ in ${\mathcal{M}}(\Lambda)$, then
    • for plan B: iterate through ${\boldsymbol{h}}^{\prime}$ in $\Lambda$ and verify that $({\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ has only been listed once.
    • for plan C: iterate through ${\boldsymbol{h}}^{\prime}$ in $\Lambda$ and verify that $({\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ has only been listed once, or else it was listed from $\sigma({\boldsymbol{h}}^{\prime})$ for some ${\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}^{\prime}}^{*}}$.
7.2. Iterative procedure for lower sets
For lower sets $\Lambda$, the cost of the above verification step can be optimized. We outline in this section how this is performed, but first introduce some necessary notation in this regard.
Given $\Lambda\subset\mathbb{N}_{0}^{d}$ lower of finite cardinality, we introduce the sets $\Lambda_{[1]},\dots,\Lambda_{[d]}$ by $\Lambda_{[j]}:=\{{\boldsymbol{h}}_{[j]}:{\boldsymbol{h}}\in\Lambda\}$. The lower property of $\Lambda$ implies that $\Lambda_{[j]}$ are also lower with $\Lambda_{[j]}\times\{0\}\subset\Lambda_{[j+1]}$ for any $j=1,\dots,d-1$. It does also imply the explicit characterization
(7.1) $\Lambda_{[j]}=\{{\boldsymbol{h}}\in\mathbb{N}_{0}^{j}:({\boldsymbol{h}},{\boldsymbol{0}})\in\Lambda\},$
where $({\boldsymbol{h}},{\boldsymbol{0}})$ means extending ${\boldsymbol{h}}$ with $d-j$ zeros. Associated mirrored sets ${\mathcal{M}}(\Lambda_{[j]})$ satisfy the same properties, i.e. ${\mathcal{M}}(\Lambda_{[j]})\times\{0\}\subset{\mathcal{M}}(\Lambda_{[j+1]})$, and ${\mathcal{M}}(\Lambda_{[j]})=\{{\boldsymbol{k}}\in\mathbb{Z}^{j}:({\boldsymbol{k}},{\boldsymbol{0}})\in{\mathcal{M}}(\Lambda)\}$. We note that
(7.2) $\Lambda_{[1]}=\{0,\dots,q\},\quad{\mathcal{M}}(\Lambda_{[1]})=\{-q,\dots,q\},\quad\quad\mbox{where}\quad\quad q=\max_{{\boldsymbol{h}}\in\Lambda}(h_{1}),$
and $\Lambda_{[d]}=\Lambda$. For the sake of convenience, we introduce the convention $\Lambda_{[0]}=\emptyset$.
For ${\boldsymbol{z}}\in\mathbb{Z}^{d}$ arbitrary and ${\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)$, there holds ${\boldsymbol{h}}\cdot{\boldsymbol{z}}={\boldsymbol{h}}_{[d-1]}\cdot{\boldsymbol{z}}_{[d-1]}$ if $h_{d}=0$ and this is equal to ${\boldsymbol{h}}_{[d-2]}\cdot{\boldsymbol{z}}_{[d-2]}$ if in addition $h_{d-1}=0$, etc. In light of this observation, we derive the lemma 7.1 below. Here again $\square$ is a placeholder for 0 (integral exactness), or A, B, or C (reconstruction using plan A, plan B or plan C).
Lemma 7.1.
Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be lower and $(n,{\boldsymbol{z}})\in\mathbb{N}\times\mathbb{Z}^{d}$. Then
  • $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{\square}(\Lambda)$ $\Longrightarrow\;(n,{\boldsymbol{z}}_{[d-1]})\in{\mathcal{P}}_{\square}(\Lambda_{[d-1]})$
    $\Longrightarrow\dots$ $\Longrightarrow\;(n,z_{1})\in{\mathcal{P}}_{\square}(\Lambda_{[1]})$.
The lemma has an immediate consequence. Namely, verifying whether $(n,{\boldsymbol{z}})$ is in ${\mathcal{P}}_{\square}(\Lambda)$ can be implemented as an iterative scheme consisting on $d$ iterations,
    $1^{st}$ iteration: check whether $(n,z_{1})$ is in ${\mathcal{P}}_{\square}(\Lambda_{[1]})$;
    $2^{nd}$ iteration: check whether $(n,{\boldsymbol{z}}_{[2]})$ is in ${\mathcal{P}}_{\square}(\Lambda_{[2]})$;
    $d^{th}$ iteration: check whether $(n,{\boldsymbol{z}})$ is in ${\mathcal{P}}_{\square}(\Lambda_{[d]})$.
The condition to be checked in each iteration can be explicitly described.
The first iteration consists in checking if $(n,z_{1})\in{\mathcal{P}}_{\square}(\{0,\dots,q\})$. For instance,
  • for $\square=0$, it consists on checking if $az_{1}\not\equiv_{n}0$ for all $a\in\{-q,\dots,q\}-\{0\}$, which is equivalent to $az_{1}\not\equiv_{n}0$ for all $a\in\{1,\dots,q\}$,
  • for $\square\in\{A,B,C\}$, it consists on checking if $(az_{1}\leavevmode\nobreak\ {\rm mod}\leavevmode\nobreak\ n)$ are pairwise distinct for $a\in\{-q,\dots,q\}$ (equivalent to $az_{1}\not\equiv_{n}0$ for $a\in\{1,\dots,2q+1\}$), but with $az_{1}\equiv_{n}-az_{1}$ allowed for $\square=C$.
It is enough to let $z_{1}=1$ and choose $n\geq q+1$, resp. $n\geq 2q+1$. In order to complete the scheme, we only need to outline how to efficiently check if $(n,{\boldsymbol{z}}_{[j]})\in{\mathcal{P}}_{\square}(\Lambda_{[j]})$ once we have established that $(n,{\boldsymbol{z}}_{[j-1]})$ is in ${\mathcal{P}}_{\square}(\Lambda_{[j-1]})$. When considering ${\boldsymbol{h}}$ in ${\mathcal{M}}(\Lambda_{[j]})$, we necessarily have ${\boldsymbol{h}}=(\widehat{\boldsymbol{h}},h_{j})$ with $\widehat{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda_{[j-1]})$ and $h_{j}=\pm p$, hence ${\boldsymbol{h}}\cdot{\boldsymbol{z}}_{[j]}=\widehat{\boldsymbol{h}}\cdot{\boldsymbol{z}}_{[j-1]}\pm pz_{j}$. For example if $\square=A$, assuming we have already computed all $\widehat{\boldsymbol{h}}\cdot{\boldsymbol{z}}_{[j-1]}$ for $\widehat{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda_{[j-1]})$ and verified they are pairwise distinct (modulo $n$), we simply iterate through ${\boldsymbol{h}}$ in ${\mathcal{M}}(\Lambda_{[j]})$ s.t. $h_{j}\neq 0$, calculate ${\boldsymbol{h}}\cdot{\boldsymbol{z}}_{[j]}$ as prescribed, and check whether the result modulo $n$ has not already been listed.
In the course of the iterative scheme, it is crucial to keep record of values ${\boldsymbol{h}}\cdot{\boldsymbol{z}}_{[j]}$ for ${\boldsymbol{h}}\in{\mathcal{M}}(\Lambda_{[j]})$ Accordingly, we introduce notation
(7.3) ${\mathcal{D}}_{j}(\Lambda):=\{V_{\boldsymbol{h}}\mbox{ for }{\boldsymbol{h}}\in\Lambda_{[j]}\},\mbox{ where }V_{\boldsymbol{h}}:=\{\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}_{[j]}:\sigma\in{\mathcal{S}}_{{\boldsymbol{h}}}\}$
Our discussion so far shows that ${\mathcal{D}}_{1}(\Lambda)\subset{\mathcal{D}}_{2}(\Lambda)\subset\dots\subset{\mathcal{D}}_{d}(\Lambda)$, and that computing ${\mathcal{D}}_{j}(\Lambda)$ is fast, with computational complexity ${\mathcal{O}}(\#({\mathcal{M}}(\Lambda_{[j]})\setminus{\mathcal{M}}(\Lambda_{[j-1]})))$. We will refer to ${\mathcal{D}}_{j}(\Lambda)$ as the dictionary associated with $\Lambda_{[j]}$. For all such dictionaries, we use notation ${\mathcal{D}}=\{{\boldsymbol{k}}\mapsto(v_{{\boldsymbol{k}}},V_{\boldsymbol{k}}^{*})\}$ refer to ${\boldsymbol{k}}$ as the key, and have $v_{{\boldsymbol{k}}}={\boldsymbol{k}}\cdot{\boldsymbol{z}}_{[j]}$ and $V_{\boldsymbol{k}}^{*}=V_{\boldsymbol{k}}-\{v_{\boldsymbol{k}}\}$. The set $\Lambda_{[j]}$ is implicit in the notation, it consists simply on all the keys. The use of dictionaries allows us to efficiently store and optimize data during the search, ensuring minimal redundant computations.
7.3. Search Algorithm
Algorithm 2 proposes an iterative procedure for obtaining $(n,{\boldsymbol{z}})$ admissible, for any given $\Lambda$ and admissibility criterion $\square\in\{0,A,B,C\}$. The algorithm initializes $n$ to a lower bound, ${\boldsymbol{z}}$ to an ‘empty vector’ and consists on $d$ iterations. Within iteration $j$, an integer $\zeta$ is incremented in $\{0,\dots,n-1\}$ until $(n,[{\boldsymbol{z}},\zeta])$ becomes admissible for $\Lambda_{[j]}$. If no suitable $\zeta$ is found, we increment $n$ and repeat the search. Once a $\zeta$ is found, ${\boldsymbol{z}}$ is extended with $\zeta$ and we advance to iteration $j+1$.
Algorithm 2 Search for admissible $(n,{\boldsymbol{z}})$
0:  $d$, $\Lambda$, $n_{\text{lower}}$, plan-$\square$.
1:  Initialize $n=n_{\text{lower}}$, ${\boldsymbol{z}}=[\leavevmode\nobreak\ ]$, ${\mathcal{D}}=\{\}$,
2:  Let $\zeta=0$, $F_{nz}=\text{True}$,
3:  for $j=1,\dots,d$ do
4:     $n\leftarrow n-1$, $\zeta\leftarrow 0$, $F_{nz}\leftarrow\text{False}$,
5:     while not $\text{not }F_{nz}$ do
6:        $n\leftarrow n+1$,
[…]
8:           $-F_{j},{\mathcal{D}}_{j}=\text{Admissibility}(\Lambda_{[j]},\leavevmode\nobreak\ {\mathcal{D}},\leavevmode\nobreak\ n,\leavevmode\nobreak\ {\boldsymbol{z}},\leavevmode\nobreak\ \zeta,\leavevmode\nobreak\ \text{plan-}\square)$- □ )
9:           if $F_{j}$ then
10:              $F_{nz}\leftarrow\text{True}$, ${\mathcal{D}}\leftarrow{\mathcal{D}}_{j}$,
11:           else
12:              $\zeta\leftarrow\zeta+1$
13:           end if
14:        end while
15:     end while
16:     ${\boldsymbol{z}}\leftarrow[{\boldsymbol{z}},\zeta]$
17:  end for
18:  return  $n$, ${\boldsymbol{z}}$, ${\mathcal{D}}$
The main procedure in every iteration is Admissibility (line 8) which evaluates the admissibility of a given pair $(n,[{\boldsymbol{z}},\zeta])$ for $\Lambda_{[j]}$, and outputs the dictionary associated with $\Lambda_{[j]}$ and $[{\boldsymbol{z}},\zeta]$ in case of admissibility. The procedure is presented in Algorithm 3.
Algorithm 3 Admissibility of $(n,[{\boldsymbol{z}},\zeta])$
[…]
1:  Initialize sets $V=\{\}$ and $V^{*}=\{\}$ and dictionary ${\mathcal{D}}_{out}=\{\}$,
[…]
3:     Extract $(v_{\text{id}},v_{\text{other}})$ associated with key $\widehat{\boldsymbol{h}}$ in ${\mathcal{D}}$,
4:     if $h\neq 0$ then
5:        update $v_{\text{id}}\leftarrow v_{\text{id}}+h\zeta$ and $v_{\text{other}}\leftarrow\{v_{\text{id}}-h\zeta\}\cup\{w\pm h\zeta\mbox{ for }w\in v_{\text{other}}\}$,
6:     end if
7:     if condition-$\square$ then
[…]
9:     else
10:        Update $V\leftarrow V\cup\{v_{\text{id}}\}$ and $V^{*}\leftarrow V^{*}\cup v_{\text{other}}$
11:        Create and associate in ${\mathcal{D}}_{out}$ the key ${\boldsymbol{h}}$ with value $(v_{\text{id}},v_{\text{other}})$.
12:     end if
13:  end for
14:  return  True, ${\mathcal{D}}_{out}$
The admissibility algorithm 3 processes the input lower set $\Lambda_{[j]}$, the dictionary ${\mathcal{D}}$ associated with $\Lambda_{[j-1]}$, the integer $n$, the vector ${\boldsymbol{z}}\in\mathbb{Z}^{j-1}$, $\zeta\in\{0,\dots,n-1\}$ and a specified admissibility for plan-$\square$ criterion. The purpose is to determine the admissibility of the couple $(n,[{\boldsymbol{z}},\zeta])$ for $\Lambda_{[j]}$ for the specified criterion. If admissibility is satisfied, the dictionary associated with $\Lambda_{[j]}$ is returned.
Admissibility for all criteria is implemented similarly. The algorithm structure is the same, except for the if condition in line 7, which varies as summarized below:
plan-$\square$ Inadmissibility Condition Time Complexity ${\mathcal{O}}(\square)$
$\square=0$ $0\in v_{\text{id}}\cup v_{\text{other}}$ ${\mathcal{O}}(\#v_{\text{other}})$
$\square=C$ $v_{\text{id}}\in V\cup V^{*}$ or $v_{\text{other}}\cap V\neq\emptyset$ ${\mathcal{O}}(\#V^{*})$
$\square=B$ $v_{\text{id}}\in v_{\text{other}}$ or $v_{\text{id}}\in V\cup V^{*}$ or $v_{\text{other}}\cap V\neq\emptyset$ ${\mathcal{O}}(\#V^{*})$
$\square=A$ repetition in $v_{\text{other}}$ or $v_{\text{id}}\in v_{\text{other}}$ or $v_{\text{id}}\in V\cup V^{*}$ or $v_{\text{other}}\cap(V\cup V^{*})\neq\emptyset$ ${\mathcal{O}}(\#V^{*})$
Table 1. Inadmissibility Conditions for each plan considered in Algorithm 3 with corresponding time complexity
Remark 7.2.
To do the time complexity analysis we make the following remarks:
  • A simple lookup has complexity of the size of the set.
  • To determine if the intersection of two lists $A$ and $B$ is non-empty, we consider the hash-set approach which has complexity ${\mathcal{O}}(\#A+\#B)$.
  • The size of $V$ is negligible compared to $V^{*}$. In fact, $V$ has the same cardinality as $\widetilde{\Lambda}$, while $V^{*}$ has same as ${\mathcal{M}}(\widetilde{\Lambda})\setminus\widetilde{\Lambda}$, where $\widetilde{\Lambda}$ consists on indices ${\boldsymbol{h}}$ that have been iterated of $\Lambda_{[j]}$.
  • $V^{*}$ are made to include all new $v_{\text{other}}$ vectors, which means #$\#v_{\text{other}}\ll\#V^{*}$ for $\Lambda$ large enough.
The conditions are meant modulo $n$. More precisely, checking if $v\in S$ means checking if $v\equiv_{n}s$ for some $s\in S$, and checking if $S\cap S^{\prime}\neq\emptyset$ means checking if $s\equiv_{n}s^{\prime}$ for some $s\in S$ and $s^{\prime}\in S^{\prime}$. Let us note that for $\square=0$, instructions in lines 1 and 10 are not needed. Also for $\square=A$, we do not need to define $V$ and $V^{*}$ separately but rather use directly $V\cup V^{*}$. When the inputs are $\Lambda_{1}=\{0,\dots,q\}$, ${\mathcal{D}}=\{\}$ and ${\boldsymbol{z}}=[\leavevmode\nobreak\ ]$, we have $\widehat{\boldsymbol{h}}=()$ is the “empty multi-index” in line 2. The convention of extracting $(v,v^{*})=(0,\emptyset)$ in line 3 and of evaluating $v\in\emptyset$ to true makes the procedure consistent with the theory.
Algorithm 2 is well posed, and finishes in finite time. For illustration, we consider plan A. We assume we are at iteration $j$, so that $(n,{\boldsymbol{z}})$ is admissible for $\Lambda_{[j-1]}$, i.e. ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda_{[j-1]})$. Then necessarily ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\neq{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda_{[j-1]})$. It is easily checked that for $\zeta$ very large, we have ${\boldsymbol{h}}\cdot[{\boldsymbol{z}},\zeta]\neq{\boldsymbol{h}}^{\prime}\cdot[{\boldsymbol{z}},\zeta]$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda_{[j]})$, and this will also holds if $\neq$ is replaced by $\not\equiv_{m}$ for $m$ also too large. As a result, with certainty we end up exiting iteration $j$ and going to iteration $j+1$.
7.4. Faster algorithm
Algorithm 2 can be very slow, in particular in high dimension. Every time $n$ is incremented and $\zeta$ is varied in $\{0,\dots,n-1\}$, new computations modulo $n$ have to be performed. We propose a new approach where finding $n$ is postponed to the end of the algorithm.
The omission of the modulo $n$ operation in the previous algorithms, particularly in the admissibility algorithm 3, results in a faster alternative. More precisely, for $\Lambda$ lower we first construct a vector ${\boldsymbol{z}}$ verifying an admissibility criterion where $\not\equiv_{n}$ is replaced by $\neq$, then after that compute an integer $n$ such that $(n,{\boldsymbol{z}})\in{\mathcal{P}}_{\square}(\Lambda)$. For example, for plan A, we first compute ${\boldsymbol{z}}$ for which holds ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\neq{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda)$, then find $n$ small such that ${\boldsymbol{h}}\cdot{\boldsymbol{z}}\not\equiv_{n}{\boldsymbol{h}}^{\prime}\cdot{\boldsymbol{z}}$ for all ${\boldsymbol{h}}\neq{\boldsymbol{h}}^{\prime}\in{\mathcal{M}}(\Lambda)$ which we can search for in {#$\{\#({\mathcal{M}}(\Lambda)),\dots,2\max_{{\boldsymbol{h}}\in{\mathcal{M}}(\Lambda)}(|{\boldsymbol{h}}\cdot{\boldsymbol{z}}|)+1\}$. While this alternative may produce a larger $n$, it significantly accelerates computations, providing a viable heuristic when prioritizing speed over optimality.
The algorithms structure is the same. We modify algorithm 2 by removing the dependence on $n$. Algorithm 4 is an iterative procedure for obtaining ${\boldsymbol{z}}$ admissible, for any given $\Lambda$ and admissibility criterion $\square\in\{0,A,B,C\}$. The algorithm initializes ${\boldsymbol{z}}$ to an ‘empty vector’ and consists on $d$ iterations. Within iteration $j$, an integer $\zeta$ is incremented in $\{0,\dots\}$ until $([{\boldsymbol{z}},\zeta])$ becomes admissible for $\Lambda_{[j]}$. Once a $\zeta$ is found, ${\boldsymbol{z}}$ is extended with $\zeta$ and we advance to iteration $j+1$. The main procedure in every iteration is Admissibility (line 6) which evaluates the admissibility of $[{\boldsymbol{z}},\zeta]$ for $\Lambda_{[j]}$, and outputs the dictionary associated with $\Lambda_{[j]}$ and $[{\boldsymbol{z}},\zeta]$ in case of admissibility. The procedure uses Algorithm 3 with the parameter $n$ being “nil” meaning that condition-$\square$ in line 7 is performed with $\neq$ instead of $\not\equiv_{n}$.
Algorithm 4 outputs ${\boldsymbol{z}}$ and associated dictionary ${\mathcal{D}}$ of all the values $\sigma({\boldsymbol{h}})\cdot{\boldsymbol{z}}$ for ${\boldsymbol{h}}\in\Lambda$. Subsequently, we can execute an algorithm that straightforwardly determines $n$ based on these outcomes as shown in Algorithm 6.
Remark 7.3.
  • For lower sets of simplex-type, such as ${\mathcal{S}}_{{\boldsymbol{1}},k}$ or in general ${\mathcal{S}}_{{\boldsymbol{w}},u}$ the above algorithms can be simplified further. For instance, for plan-A admissibility, it is computationally more advantageous to rely (6.5) instead of (3.21) while for plan-C admissibility, it is computationally more advantageous to rely (6.6) instead of (3.23).
  • Non optimality of the idea of Algorithm 4 + Algorithm 6 should be discussed.
Algorithm 4 Search for admissible ${\boldsymbol{z}}$
0:  $d$, $\Lambda$, plan-$\square$.
1:  Initialize ${\boldsymbol{z}}=[\leavevmode\nobreak\ ]$, ${\mathcal{D}}=\{\}$,
2:  Let $\zeta=0$, $F=\text{True}$
3:  for $j=1,\dots,d$ do
4:     $\zeta\leftarrow 0$, $F\leftarrow\text{False}$,
5:     while not $\text{not }F$ do
6:        $-F_{j},{\mathcal{D}}_{j}=\text{Admissibility}(\Lambda_{j},\leavevmode\nobreak\ {\mathcal{D}},\leavevmode\nobreak\ nil,\leavevmode\nobreak\ {\boldsymbol{z}},\leavevmode\nobreak\ \zeta,\leavevmode\nobreak\ \text{plan-}\square)$- □ )
7:        if $F_{j}$ then
8:           $F\leftarrow True$, ${\mathcal{D}}\leftarrow{\mathcal{D}}_{j}$,
9:        else
10:           $\zeta\leftarrow\zeta+1$
11:        end if
12:     end while
13:     ${\boldsymbol{z}}\leftarrow[{\boldsymbol{z}},\zeta]$
14:  end for
15:  return  ${\boldsymbol{z}}$, ${\mathcal{D}}$
Algorithm 5 admit_n
0:  $\Lambda$, ${\mathcal{D}}$, $n$, plan-$\square$
1:  Let $V=\{\}$ and $V^{*}=\{\}$,
2:  for each ${\boldsymbol{h}}$ in $\Lambda$ do
3:     Extract $(v_{\text{id}},v_{\text{other}})$ from ${\mathcal{D}}$, associated with key ${\boldsymbol{h}}$,
4:     if condition-$\square$ modulo $n$ then
5:        return  False,
6:     else
7:        Update $V\leftarrow V\cup\{v_{\text{id}}\}$ and $V^{*}\leftarrow V^{*}\cup\{v_{\text{other}}\}$
8:     end if
9:  end for
10:  return  True
Algorithm 6 Search for admissible $n$ given ${\boldsymbol{z}}$
0:  $\Lambda$, ${\mathcal{D}}$, $n_{\text{lower}}$, plan-$\square$.
1:  Initialize $n=n_{\text{lower}}-1$, $F=\text{False}$,
2:  while not $\text{not }F$ do
3:     $n\leftarrow n+1$,
4:     $-F=\text{admit\_n}(\Lambda,\leavevmode\nobreak\ {\mathcal{D}},\leavevmode\nobreak\ n,\leavevmode\nobreak\ \text{plan-}\square)$- □ ) {admit_n is defined in Algorithm 5}
5:  end while
6:  return  $n$
8. Performance Analysis
8.1. Complexity Analysis
  • Algorithm 3: The outer loop iterates over each element in $\Lambda_{[j]}$. Updating $v_{\text{other}}$ is done in $\mathcal{O}(2^{|{{\boldsymbol{h}}}|_{0}})$ time. The check of condition-$\square$ has complexity $\mathcal{O}(\#V^{*})$ negligeable with respect to ${\mathcal{O}}(\#\mathcal{M}(\Lambda_{[j]}))$ as per Table 1. Thus, the overall complexity of Algorithm 3 can be expressed at worst as
    $\mathcal{O}(\#\Lambda_{[j]}\times\#\mathcal{M}(\Lambda_{[j]}))$
  • Algorithm 2: The outer loop runs $d$ times, corresponding to each dimension. The inner WHILE loop continues until a valid $\zeta$ is found, with a maximum of $n_{j}$ iterations. For each iteration, the ”Admissibility” function is called with complexity $\mathcal{O}(\#\Lambda_{[j]}\times\#\mathcal{M}(\Lambda_{[j]}))$. Therefore, the overall worst-case complexity is
    $\mathcal{O}(p_{\square}^{*}(\Lambda)^{2}\times\#\Lambda\times\#\mathcal{M}(\Lambda)).$
  • Algorithm 4: Since this algorithm does not depend on $n$, its complexity is reduced to
    $\mathcal{O}(\#\Lambda\times\#\mathcal{M}(\Lambda)).$
  • Algorithm 5: The outer FOR loop runs once for each element ${\boldsymbol{h}}$ in $\Lambda$, making the loop run #$\#\Lambda$ times. Checking condition-$\square$ modulo $n$ has complexity $\mathcal{O}(\square)$(Table 1). Thus, following the same logic as per Algorithm 3, the total time complexity is
    $\mathcal{O}(\#\Lambda\times\#\mathcal{M}(\Lambda)).$
  • Algorithm 6: The algorithm iterates over $n$ using Algorithm 5 until it finds a good $n$ with worst case being $p_{\square}^{*}(\Lambda)$, achieving a worst-case time complexity of:
    $\mathcal{O}(p_{\square}^{*}(\Lambda)\times\#\Lambda\times\#\mathcal{M}(\Lambda)).$
8.2. Experimental Results
8.2.1. Comparison to minimal values
For our numerical investigation, we utilize anisotropic simplex-type sets in three dimensions, characterized by ${\boldsymbol{w}}=(0.9,0.8,0.7)$. Instead of specifying a radius $u$, we parametrized the sets ${\mathcal{S}}_{{\boldsymbol{w}},u}$ by their cardinalities. More precisely, given an integer $N\geq 1$, we generate $\Lambda\subset\mathbb{N}^{d}$ consisting in the $N$ multi-indices ${\boldsymbol{h}}$ with the smallest values of $\langle{\boldsymbol{h}},{\boldsymbol{w}}\rangle$ with ties broken arbitrarily. For give sizes of $\Lambda$, we employ an Algorithm 1 to compute the optimal $n_{\square}^{*}(\Lambda)$ for each plan-$\square$, and compare to the resulting values obtained using our approach (Algorithm 4 - 5). Table 2 illustrates the diverse values obtained. Our algorithm yields identical results across all three plans.
#$\#(\Lambda)$$\#({\mathcal{M}}(\Lambda))$ $n_{A}^{*}(\Lambda)$ $n_{B}^{*}(\Lambda)$ $n_{C}^{*}(\Lambda)$ $n_{A}(\Lambda)=n_{B}(\Lambda)=n_{C}(\Lambda)$
40 149 160 151 151 180
50 195 217 200 200 253
80 341 379 363 362 418
100 445 481 471 471 591
150 713 803 759 759 962
Table 2. Comparison of optimal $n_{\square}^{*}(\Lambda)$ and values obtained using our approach for different sizes of $\Lambda$.
We compute the Mean Absolute Percentage Error (MAPE) given by the formula:
$\text{MAPE}_{\square}=\frac{100\%}{N}\sum_{i=1}^{N}\left|\frac{n_{\square}^{*}(\Lambda_{i})-n_{\square}(\Lambda_{i})}{n_{\square}^{*}(\Lambda_{i})}\right|.$
This metric expresses the error as a percentage of the actual values, which can be more intuitive for understanding relative accuracy.
$\text{MAPE}_{A}$ $\text{MAPE}_{B}$ $\text{MAPE}_{C}$
16.42% 17.78% 18.3%
Table 3. MAPE for the different plans
In Table 3, we observe an overall deviation of around $17\%$, suggesting that our algorithm produces reasonably accurate estimates swiftly. While this margin of error may appear substantial, it serves as a useful starting point for search algorithms or offers quick insights into the necessary number of points.
We only consider dimension $d=3$ due to the increased complexity and time requirements associated with reporting the greedy search in higher dimensions ${\mathcal{O}}(n^{*}_{\square}(\Lambda)^{d+1})$.
8.2.2. Execution Time
For the second setup, we consider ${\boldsymbol{w}}_{3}=(0.9,0.8,0.7)$, ${\boldsymbol{w}}_{4}=(0.9,0.8,0.7,0.6)$, ${\boldsymbol{w}}_{5}=(0.9,0.8,0.7,0.6,0.5)$, and ${\boldsymbol{w}}_{8}=(0.9,0.8,0.7,0.6,0.5,0.4,0.3,0.2)$ for the cases $d=3,4,5$ and $d=8$ respectively. We consider for each dimension sets $\Lambda$ of different cardinalities. The scatter plot in Figure 2 illustrates the execution time for the different plans (Plan A, Plan B, and Plan C) across the dimensions for different choices of #$\#(\Lambda)$ using Algorithms4-6.
[…]
A linear regression analysis yields $p\lesssim 1$ for different dimensions, supporting our complexity analysis.
Finally, we make the following observations:
  • Dimension has minimal impact on our algorithm, as the effect of dimension is encapsulated within ${\mathcal{M}}(\Lambda)$. For the same #$\#(\Lambda)$, #$\#({\mathcal{M}}(\Lambda))$ increases with dimension.
  • The overall complexity can be bounded using the inequality #$\#{\mathcal{M}}(\Lambda)\leq\min\{2^{d}\#\Lambda,\#\Lambda^{\ln 3/\ln 2}\}$ from [ChkifaLower]*Lemma 3.3. Thus, our algorithm’s complexity is bounded by
    ${\mathcal{O}}(p_{\square}^{*}(\Lambda)\times\min\{2^{d}\#\Lambda^{2},\#\Lambda^{\ln 3/\ln 2+1}\}).$
  • The execution time values provided are indicative, obtained from computations on a machine with an Intel i7 CPU, 31.2 GB of RAM, and a 233.18 GB disk. All our algorithms are single-core implementations..
9. Conclusion
In this paper, we have addressed the challenge of constructing optimal rank-1 lattices for function integration and reconstruction, specifically focusing on finite index sets that are lower sets namely blocks and simplexes. We introduced new theoretical sharp lower bounds on the minimal number of nodes needed for reconstruction and developed a novel algorithm for constructing the corresponding optimal generator vector of the rank-1 lattice. Our approach, which leverages the inherent lower structure of the set of interpolants, optimizes the construction process through exhaustive search and meticulous verification, resulting in significant improvements in computation time and memory usage. The efficiency and practicality of our algorithm were demonstrated through numerical experiments, confirming its suitability for handling large-scale problems. Future research may explore further refinements of the algorithm and its application to other functional spaces and integration tasks.
Figure 2. Execution time patterns across the three plans for different dimensions as function of #$\#\Lambda$ (left) and #$\#{\mathcal{M}}(\Lambda)$ (middle) and #Λ#($\#\Lambda\#{\mathcal{M}}(\Lambda)$ (right).